### v1

# Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy Lensing

Abstract
[…]
keywords:
gravitational lensing: weak – galaxies: haloes – cosmology: observations – (cosmology:) cosmic background radiation
pubyear: 2015pagerange: Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy LensingB
1 Introduction
With large imaging surveys like the Dark Energy Survey (DES) (Abbott et al., 2022), the Kilo-Degree Survey (KiDS) (Heymans et al., 2021), and the Hyper Suprime-Cam Survey (HSC) (Sugiyama et al., 2023), we have seen a rapid change in our ability to probe the large scale structure (LSS) and matter distribution in the non-linear regime using weak lensing (WL) and clustering. One of the biggest challenges standing in the way of using such measurements to test the dark components of the standard model is the modelling of baryons. In particular, baryonic feedback suppresses the matter power spectrum on non-linear scales, and therefore the weak lensing amplitude, and must be modelled accurately to avoid biasing cosmological inference (e.g., van Daalen et al., 2011a; Chisari et al., 2019a; Bigwood et al., 2024).
However, the complexity of baryonic feedback requires a large dynamic range in resolution and time that is difficult to simulate (e.g. Crain & van de Voort, 2023). This has led to the development of simplified ‘subgrid’ prescription models in hydrodynamical cosmological simulations which are used to mimic the effects of baryonic feedback on LSS (Steinborn et al., 2015; Chisari et al., 2019a; Schaye et al., 2023; Pakmor et al., 2023). Although the simulations are able to reproduce a range of observations, they give a wide spread of predictions for the amplitude and scale extent of the suppression of the matter power spectrum due to baryonic feedback (e.g., Chisari et al., 2019a; Villaescusa-Navarro et al., 2021; Sunseri et al., 2023). Understanding and reducing this model uncertainty is the challenge. More recently, it has been possible to constrain the matter power suppression from a number of observables (e.g. Debackere et al., 2020; Grandis et al., 2024; Bigwood et al., 2024; La Posta et al., 2024; Medlock et al., 2025), although these constraints are still not very precise.
For this reason, WL and clustering cosmological analyses lose a substantial fraction of their constraining power to the uncertainty in modelling baryon feedback (see, for example, Fig. 12 of Amon et al., 2022). Many analyses take the approach of throwing away measurements on small angular scales to circumvent modelling altogether (e.g. Prat et al., 2022; Zacharegkas et al., 2022; Amon et al., 2023; Lange et al., 2023). Even when attempting to model feedback, the uncertainty and uninformed parameter space is costly (e.g. Huang et al., 2019; Schneider et al., 2022; Aricò et al., 2023; Bigwood et al., 2024), unless informed by simulations (Asgari et al., 2021; Bigwood et al., 2024; Zennaro et al., 2024). Another approach is to jointly model the WL and clustering data with probes of the baryon content (Debackere et al., 2020; Tröster et al., 2022; Schneider et al., 2022; Bigwood et al., 2024).
The kinetic Sunyaev-Zeldovich (kSZ) effect is a secondary anisotropy in the CMB that comes from the Doppler boost of CMB photons by inverse Compton scattering of free electrons in the Intracluster Medium (ICM) and Circumgalactic Medium (CGM) that have bulk motion along the line of sight relative to the CMB rest frame (Zeldovich & Sunyaev, 1969; Mroczkowski et al., 2019). The kSZ effect is small relative to other CMB secondaries (e.g. tSZ), follows a blackbody spectrum, and is directly proportional to the product of the density of electrons surrounding galaxies and their bulk velocity along the line of sight. The dependency on the bulk line of sight velocity and the small amplitude relative to other secondaries requires methods such as a velocity weighted stack using external velocity measurements from galaxy surveys to measure the signal. Stacked measurements with a velocity reconstruction weighting have revealed that the electrons around galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS) are significantly more extended than the dark matter distribution (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a). The kSZ effect is particularly useful for WL as it probes the density of electrons (and therefore baryons) in the outskirts of the halo ($\sim$1-4 virial radii) which lies in the range of scales that baryons need to be better understood to improve WL cosmological constraints. Joint analyses of the kSZ effect with WL measurements are a promising way to constrain baryonic feedback and simultaneously improve cosmological constraints (e.g., Bigwood et al., 2024).
In order to accurately interpret the kSZ effect, one must know the halo mass of the sample (Moser et al., 2021) and have an appropriate forward model as was done in Amodeo et al. (2021). By jointly analysing the kSZ effect with GGL measurements we can build a self-consistent framework for estimating the halo mass. Without GGL, analyses of the kSZ effect either rely on estimates of stellar mass (Maraston & Strömbäck, 2011; Maraston et al., 2013) which are then used to estimate the halo mass through stellar-to-halo mass relations (Miyatake et al., 2015; Kravtsov et al., 2018), or fits to galaxy clustering (Miyatake et al., 2015; Rodríguez-Torres et al., 2016), both of which have greater uncertainties than galaxy-galaxy lensing estimates. Bigwood et al. (2024) estimated that for the BOSS CMASS galaxy sample, these estimates span almost an order of magnitude. In this work we present the first joint fits of kSZ measurements and GGL measurements to disentangle baryons from dark matter in the outskirts of galactic halos, and more accurately constrain the mass of the sample. Accurate mass estimates where the baryon content is considered is also important for cluster cosmology (e.g. Cromer et al., 2022).
Measurements of the kSZ effect can be used to test baryon feedback models and hydrodynamical simulations. In this case, having an accurate halo mass of the sample is as important, in order to interpret the appropriate strength of feedback (McCarthy et al., 2024). Previous analyses of the BOSS kSZ measurements of Schaan et al. (2021) are consistent with feedback that is stronger than most hydrodynamical simulations predict (Bigwood et al., 2024; McCarthy et al., 2024). This underestimation may be severe enough that it may even bias WL cosmological constraints (Amon & Efstathiou, 2022; Preston et al., 2023). In this work we build a fast analytic modelling framework based on the halo model, complementary to previous approaches that rely on the baryonification method, or expensive hydrodynamical simulations. This framework can be used for cosmological analyses of GGL and clustering in the future.
This work is structured as follows: Section 2 describes the kSZ and GGL measurements that we jointly analyse, Section 3 details the model that we construct, Section 4 demonstrates the model in practice and presents the joint fits for BOSS galaxies and Section 5 concludes the work.
2 kSZ & galaxy-galaxy lensing measurements
In this work, we build a model to consistently describe the excess surface mass density of halos, as measured by GGL (Amon et al., 2023), with the gas distribution, probed by measurements of the kSZ effect (Schaan et al., 2021). We describe these measurements in more detail below.
2.1 Galaxy-Galaxy Lensing
In this work we use GGL measurements from Amon et al. (2023) [^1] https://github.com/aamon/Galaxy-galaxy-lensing-KiDS-DES-HSC to determine the excess surface mass density profile of BOSS DR12 CMASS galaxies. These are obtained from a cross-correlation of the spectroscopically determined positions of the BOSS galaxies with the photometric shapes of background source galaxies of the Dark Energy Survey Year 3 (DES Y3) and the Kilo-Degree Survey 1000 square degree (KiDS-1000, 4th KiDS Data Release) data.
The DES Y3 data were collected with the Dark Energy Camera on the 4-meter Blanco Telescope and spans 4143 $\rm deg^{2}$ of 5-band photometry (grizY) with an average number density of 5.59 arcmin-2. The data include shapes of over 100 million galaxies across four tomographic redshift bins (Myles et al., 2021). The shapes that enabled these measurements were presented in Gatti et al. (2021).
The KiDS-1000 weak lensing data were taken with the 2.6 m VLT Survey Telescope (VST) (Capaccioli & Schipani, 2011; Grado et al., 2012) and span 777 $\rm deg^{2}$ of the sky (Giblin et al., 2021) with a number density of 6.22 $\rm arcmin^{-2}$.
Given spectroscopically measured redshifts of lenses and photometric shape and redshift measurements of the sources, the quantity that is directly measured is the average tangential shear $\langle\gamma_{t}\rangle$ as a function of angular separation $\theta$. We can then use this to determine the differential surface mass density $\Delta\Sigma$
$\Delta\Sigma(R)=\frac{\langle\gamma_{t}(\theta)\rangle}{\overline{\Sigma_{\rm crit}^{-1}}}\;,$ (1)
where the angular separation is related to the projected distance by $\theta=R/\chi(z_{l})$. The denominator is referred to as the average inverse critical surface mass density, and is a function of the redshifts of both the lens $z_{l}$ and the source $z_{s}$. For an individual lens-source pair, $\Sigma_{\rm crit}^{-1}$ is
$\Sigma_{\rm crit}^{-1}(z_{l},z_{s})=\frac{4\pi G}{c^{2}}\frac{\chi(z_{l})[\chi(z_{l})-\chi(z_{s})]}{\chi(z_{s})a(z_{l})}\;,$ (2)
where $\chi$ is the line-of-sight comoving distance. In practice we measure $\Delta\Sigma$ using an estimator of the form
$\Delta\Sigma=\frac{\sum_{ls}w_{\mathrm{sys},l}w_{ls}\Sigma_{\mathrm{crit}}(z_{l},z_{s})e_{t}}{\sum_{ls}w_{\mathrm{sys},l}w_{ls}}\,.$ (3)
where $\sum_{ls}$ refers to the sum of all lens-source pairs, $w_{\mathrm{sys},l}$ is the systematic weight of each lens provided by the lens survey to prevent selection bias, $e_{t}$ is the tangential ellipticity of a source, and $w_{s}$ is a source weight provided from the source survey to minimize systematic errors in the survey which we use to create a combined lens-source weight $w_{\rm ls}$,
$w_{\rm ls}=w_{s}\left(\Sigma_{\mathrm{crit}}^{-1}(z_{l},z_{s})\right)^{2}\;.$ (4)
This is a basic estimator, which is calibrated with additional corrections to account for systematics such as photometric redshift uncertainties, shear response and calibration, overlapping lens and source populations (boost factors) and spurious systematic signals (random corrections). We refer to Amon et al. (2023) for a detailed description of all the relevant systematic corrections that are required to yield a robust measurement of $\Delta\Sigma$.
2.2 kSZ Effect
The kinetic SZ effect refers to the contribution to the SZ effect that comes from the bulk motion of the electrons in galactic halos. For this reason, the kSZ effect is directly proportional to the number density of electrons and their bulk velocity along the line of sight (LoS). The temperature fluctuations from the kSZ effect can be written as
$\frac{\delta T_{\rm kSZ}}{T_{\rm CMB}}=\frac{\sigma_{T}}{c}\int_{\rm LoS}\frac{d\chi}{1+z}n_{\rm e}(\chi)v_{\rm p}e^{-\tau}\;,$ (5)
where $\sigma_{T}$ is the Thompson cross-section, $c$ is the speed of light, $\chi$ is the comoving distance to redshift $z$, $n_{\rm e}$ is the physical (not comoving) electron number density, $v_{\rm p}$ is the peculiar velocity along the line of sight, and $\tau$ is the optical depth to Thompson scattering.
The kSZ effect measurements used in this work were obtained by Schaan et al. (2021) by stacking the CMB temperature fluctuations mapped by ACT (in 98 GHz and 150 GHz frequencies) around BOSS CMASS galaxies (DR10), weighted by their peculiar velocities. The covariance matrix is constructed using the bootstrap method. [^2] Higher signal-to-noise kSZ measurements have recently been reported in (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a) using ACT DR6 data; our method could be extended in future to incorporate these data.
The coadded CMB temperature maps were created by combining the ACT DR5 day & night maps (Fowler et al., 2007; Swetz et al., 2011; Thornton et al., 2016; Henderson et al., 2016) with data from the Planck satellite (Planck Collaboration et al., 2020a). The beams used to create the f90 and f150 ACT DR5 maps are nearly Gaussian with FWHM of 2.1 and 1.3 arcminutes respectively (Lungu et al., 2022). The kSZ effect measurements of Schaan et al. (2021) can be extracted from a temperature map $\delta T$ by using compensated aperture photometry (CAP) filtering with varying aperture radius $\theta_{\rm d}$. The CAP filter outputs an integrated temperature $\mathcal{T}$ in the disk with radius $\theta_{\rm d}$ subtracted by the temperature in a concentric ring of the same area around the original disk. Mathematically this is written as
$\mathcal{T}(\theta_{\rm d})=\int d^{2}\theta\delta T(\theta)W_{\theta_{\rm d}}(\theta)\;,$ (6)
where the CAP filter $W_{\theta_{\rm d}}(\theta)$ is defined as
$W_{\theta_{\rm d}}(\theta)=\begin{cases}1&\theta<\theta_{\rm d}\\ -1&\theta_{\rm d}\leq\theta\leq\sqrt{2}\theta_{\rm d}\\ 0&\text{else}\end{cases}\;.$ (7)
The measurements span a range of aperture radii $\theta_{\rm d}$ from 1 to 6 arcminutes which probes the kSZ profile in the range of 0.5 to 4 virial radii in the CMASS galaxy sample. These scales were chosen to probe baryonic feedback while minimizing covariance in the measurements. Schaan et al. (2021) stacked the measured temperatures $\mathcal{T}_{i}(\theta_{\rm d})$ around each galaxy in the sample. The stacking is done using a velocity weighted, inverse-variance weighted mean estimator given by
$\hat{T}_{\rm kSZ}(\theta_{d})=-\frac{1}{r_{v}}\frac{v^{\rm rec}_{\rm rms}}{c}\frac{\sum_{i}\mathcal{T}_{i}(\theta_{d})\left(\frac{v^{\rm rec}_{i}}{c}\right)\frac{1}{\sigma_{i}^{2}}}{\sum_{i}\left(\frac{v^{\rm rec}_{i}}{c}\right)\frac{1}{\sigma_{i}^{2}}}\;.$ (8)
We define $\sigma_{i}$ as the noise on the CAP filter for the $i$th galaxy. We note that the bias factor $r_{v}$, RMS reconstructed radial velocity $v^{\rm rec}_{\rm rms}$, and the reconstructed radial velocity of each galaxy $v^{\rm rec}_{i}$ are defined from the velocity reconstruction procedure outlined in Schaan et al. (2021) which is similar to the displacement reconstruction done in Baryon Acoustic Oscillation (BAO) studies (Padmanabhan et al., 2012). The velocity reconstruction procedure involves solving the linearized continuity equation in redshift-space
$\mathbf{\nabla}\cdot\mathbf{v}+f\mathbf{\nabla}\cdot[(\mathbf{v}\cdot\mathbf{\hat{n}})\mathbf{\hat{n}}]=-aHf\frac{\delta_{\rm g}}{b}\;,$ (9)
where $\mathbf{v}$ is the velocity field, $f=d\ln\delta/d\ln a$ is the logarithmic linear growth rate, $\mathbf{\hat{n}}$ is the line-of-sight, $a$ is the scale factor, $H$ is the Hubble parameter, $b$ is the linear bias, and $\delta_{\rm g}$ is the galaxy overdensity. Solving the linearized continuity equation allows for the reconstruction of the radial velocity field which can be assigned to each galaxy in the sample. The estimator is carefully chosen so that the radial direction of the velocity field does not cancel out thus leaving the kSZ signal intact. Lastly, the measurements of Schaan et al. (2021) perform a bootstrap resampling of the individual galaxies to estimate the covariance of the stacked kSZ profiles. The measurements at smaller aperture sizes have a dominant source of noise coming from the detector and foregrounds while the measurements at larger aperture sizes are dominated by large-scale CMB fluctuations.
3 Halo model
To model the kSZ and GGL signals we model the density profiles of both baryons and dark matter around our lens galaxy sample. In this work we start by assuming the density profiles are axisymmetric. We can separate the matter density profile into two components
$\rho_{\rm m}(r)=\rho_{\rm b}(r)+\rho_{\rm dm}(r)\;,$ (10)
where $\rho_{\rm b}(r)$ is the baryon density profile and $\rho_{\rm dm}(r)$ is the dark matter density profile. In this work, we are modelling the density profiles of matter at scales of order the virial radius and thus in our model we assume that the baryons are fully described by the gas such that $\rho_{\rm b}(r)\approx\rho_{\rm gas}(r)$ where $\rho_{\rm gas}(r)$ is the gas density profile. We can expect that the fraction of baryons in stars and cold non-ionized gas is $\lesssim 10\%$ (Kelly et al., 2021), and since baryons make up $\sim 15\%$ of the total matter, we estimate that this assumption should not lead to an error larger than $\sim 1.5\%$ in the total matter density. Each density profile in our model has two components
$\rho(r)=\rho^{\rm 1h}(r)+\rho^{\rm 2h}(r)\;.$ (11)
The first term refers to the 1-halo density profile which corresponds to density from the effective central halo, and the second term refers to the 2-halo density profile which corresponds to contributions in the density coming from the surrounding neighbouring halos (Asgari et al., 2023). We model the transition between the 1-halo and 2-halo terms as a sum. This simplistic treatment of the transition region should suffice for the current level of precision in our analysis. In the subsequent subsections below we will explain these 1-halo and 2-halo terms in more detail.
We use a fixed Planck flat $\Lambda$CDM cosmology with the following values for parameters: $\Omega_{\rm m}=0.3111$, $\Omega_{\rm b}=0.0490$, $h=0.6766$, $n_{s}=0.9665$, $\sigma_{8}=0.8102$ (Planck Collaboration et al., 2020b). We fix the redshift of our model to be $\langle z\rangle=0.55$, the mean redshift of our lens sample (BOSS CMASS). This choice does not impact the results of our analysis when compared to a marginalization over the redshift distribution of the lens sample as shown in Amodeo et al. (2021). We further verified this to be true while including GGL.
3.1 The One Halo Term
The 1-halo term of the density profile is the contribution to the density from the central halo itself. For this work we assume the host galaxy lies at the centre of a dark matter halo and that the gas is bound to the dark matter halo (i.e. no miscentering). In other words, we do not model the gas and dark matter around satellites and centrals with a Halo-Occupation Distribution (HOD) framework. This model ignores the potential systematic effect of satellites or ‘miscentering’, where satellite galaxies offset the stacking procedures used on the central galaxies in both the kSZ effect measurements and the GGL measurements.
We also ignore an additional uncertainty on the reconstruction bias factor $r_{v}$ which is about $\sim 10\%$ for the reconstruction used in this analysis (Schaan et al., 2021). We note that modelling of more precise measurements should include this effect (Ried Guachalla et al., 2024; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a).
3.1.1 Dark Matter
We model the 1-halo dark matter density profile as a Navarro-Frank-White (NFW) profile Navarro et al. (1997) which follows the form
$\rho_{\rm NFW}(r)=\frac{\delta_{\rm c}\bar{\rho}_{\rm m}}{\left(\frac{r}{r_{\rm s}}\right)\left(1+\frac{r}{r_{\rm s}}\right)^{2}}\;,$ (12)
where the scale radius $r_{\rm s}$ is related to the comoving spherical overdensity halo radius $r_{\Delta_{\rm m}}$ by the concentration parameter $c(M)$ by the relation $r_{\Delta_{\rm m}}=c(M,z)r_{\rm s}$ and $\bar{\rho}_{\rm m}$ is the comoving average cosmic matter density. We define the characteristic halo overdensity $\delta_{\rm c}$ as
$\delta_{\rm c}=\frac{\Delta c^{3}}{3[\log(1+c)-c/(1+c)]}\;,$ (13)
where $\Delta$ is the spherical overdensity of the halo determining the mass definition of the halo. The comoving spherical overdensity halo radius $r_{\Delta}$ can be written as
$r_{\Delta_{\rm m}}=\left(\frac{3M_{\rm halo}}{4\pi\Delta\bar{\rho}_{\rm m}}\right)^{1/3}\;.$ (14)
In our work we set the spherical overdensity $\Delta=200$ and use a concentration-mass relation given by Duffy et al. (2008)
$c(M_{\rm halo},z)=5.71\times\left(\frac{M_{\rm halo}}{2\times 10^{12}\;\rm M_{\odot}}\right)^{-0.084}(1+z)^{-0.47}\;,$ (15)
to determine the concentration given a halo mass $c(M_{\rm halo})$ at $z=0.55$. By using $\Delta=200$, our definition of halo mass $M_{\rm halo}$ is the same as $M_{200}$. This concentration relation was determined from N-body simulations for halo masses $10^{11}-10^{15}\;\rm M_{\odot}$ and redshifts $20<z<$ 2 which matches our selected sample. We use a comoving $\bar{\rho}_{\rm m}$ because the radius $r$ is in comoving units. With this
$\rho^{\rm 1h}_{\rm dm}(r)=f_{\rm dm}\rho_{\rm NFW}(r)\;,$ (16)
where $f_{\rm dm}$ is the universal dark matter fraction of the universe defined as $f_{\rm dm}\equiv\Omega_{\rm dm}/\Omega_{\rm m}$. In our model, the matter in the universe is only composed of dark matter and baryons, so $f_{\rm dm}=1-f_{\rm b}$ where $f_{\rm b}$ is the universal baryon fraction defined as $f_{\rm b}\equiv\Omega_{\rm b}/\Omega_{\rm m}$.
In this work, we don’t model the potential effects of assembly bias from the surrounding environment in our CMASS sample. This effect could bias our results at the few percent level (Yuan et al., 2021; Amon et al., 2023) and is a necessary improvement for future work. We also choose to ignore the back-reaction of baryonic feedback on the dark matter distribution: a second-order effect which is not necessary to model for our current level of precision (van Daalen et al., 2011b; Springel et al., 2018). Our model of the 1-halo term closely follows an approach from Cromer et al. (2022) which modelled baryons in WL of clusters. This approach differs from HOD based halo models commonly used to model GGL (Nishimichi et al., 2019; Miyatake et al., 2022; Zacharegkas et al., 2022; Dvornik et al., 2023), but it does allow for an explicit modelling of the baryon density profile which can be used to jointly analyse the kSZ measurements enabling constraints on baryon density parameters and dark matter parameters simultaneously.
Figure 1: The impact of model parameters on the quantities of interest: baryon density profile (left), kSZ temperature profile at 150 GHz (middle), and excess surface mass density $\Delta\Sigma$ (right). The model contains 6 parameters varied from top to bottom: halo mass $M_{\rm halo}$, core scale fraction $x_{\rm c}$, intermediate power law slope $\alpha$, outer power law slope $\beta$, inner power law slope $\gamma$, and the 2-halo amplitude factor $A_{\rm 2h}$. For our fiducial model we hold several parameters fixed to alleviate degeneracy between parameters. In the fiducial model we hold $\alpha=1$ and $\gamma=0.2$ fixed while using a concentration mass relation $c(M_{\rm halo})$. In the left most panel we show where $x=1$ ($r=r_{\rm 200c}$). We found the halo mass $M_{\rm halo}$, the outer power law slope $\beta$, and the 2-halo amplitude $A_{\rm 2h}$ to be the most impactful parameters in the model given the scales we are probing with the kSZ effect and GGL. There are degeneracies between these parameters when only one measurement (kSZ or GGL) is being fit to which can be seen in either the $T_{\rm kSZ}$ or $R\Delta\Sigma$ panels.
kSZ Only GGL Only kSZ + GGL
Parameter Description Prior 68% limits 68% limits 68% limits
$\log_{10}(M_{\rm halo})$ Halo Mass $\mathcal{U}(12.5,14.0)$ $13.42^{+0.44}_{-0.27}$ $13.41^{+0.08}_{-0.07}$ $13.44^{+0.06}_{-0.05}$
$x_{c}$ Core scale fraction $\mathcal{U}(0.0,1.0)$ $>0.48$ $>0.46$ $>0.49$
$\alpha$ Intermediate slope
$\beta$ Outer slope $\mathcal{U}(1,5)$ $2.47^{+0.34}_{-0.46}$ $<3.51$ $2.45^{+0.30}_{-0.23}$
$\gamma$ Inner slope
$A_{\rm 2h}$ Two halo term amplitude $\mathcal{U}(0.0,1.5)$ $>0.631$ $0.66\pm 0.16$ $0.64\pm 0.16$
Table 1: Summary of the halo model parameters, their priors, and the 68% confidence intervals for the fit to only kSZ measurements, only GGL measurements, and joint kSZ and GGL data. The bold font indicates the parameters that are allowed to vary in the fit. In our fiducial model, we use a concentration-mass relation and we fix $\alpha=1.0$ and $\gamma=0.2$ to alleviate degeneracies in the GNFW model.
3.1.2 Baryons
We model the 1-halo density profile of baryons as a single Generalized NFW (GNFW) profile described in Zhao (1996); Battaglia (2016); Amodeo et al. (2021); Cromer et al. (2022) which follows the form
$\rho_{\rm GNFW}(r)=\rho_{0}\left(\frac{x}{x_{\rm c}}\right)^{-\gamma}\left(1+\left(\frac{x}{x_{\rm c}}\right)^{1/\alpha}\right)^{-(\beta-\gamma)\alpha}\;,$ (17)
This profile is described by five parameters: $\rho_{0}$ the density amplitude of the gas, $\alpha$ the central power law slope ($x\sim 1$), $\beta$ the outer power law slope ($x\gg 1$), $\gamma$ the inner power law slope ($x\ll 1$), and $x_{\rm c}$ the core scale fraction. [^3] We note that there are several slightly different versions of this profile in the literature, this implementation follows the original proposed definition of Zhao (1996) where $\alpha,\beta,\gamma=1,3,1$ corresponds to a standard NFW profile power law scaling. We define
$x\equiv\frac{r}{r_{\rm\Delta_{\rm c}}}=r\left(\frac{4\pi\Delta\rho_{\rm crit}}{3M_{\rm halo}}\right)^{1/3}\;,$ (18)
where we again use a spherical overdensity of $\Delta=200$, but now we use the comoving critical density of the universe $\rho_{\rm crit}$ instead of the comoving average matter density $\bar{\rho}_{\rm m}$. We use a comoving $\rho_{\rm crit}$ because the radius $r$ is in comoving units. From here the 1-halo term of the baryon density profile is given by
$\rho^{\rm 1h}_{\rm b}(r)=f_{\rm b}\rho_{\rm GNFW}(r)\;.$ (19)
where $f_{\rm b}$ is the universal baryon fraction of the universe defined as $f_{\rm b}\equiv\Omega_{\rm b}/\Omega_{\rm m}$. It is important to note that $\rho_{0}$ is responsible for setting the amplitude of the baryon density profile, therefore we must be careful to ensure that the amplitude of the baryon density profile is normalized such that at sufficiently large scales (several virial radii) the mass fraction of baryons reaches the cosmic baryon fraction $f_{\rm b}$. This normalization can only be done numerically and is described in Cromer et al. (2022). We choose $\rho_{0}$ such that the mass ratio of baryons to total matter enclosed in a sphere of radius $r_{\rm b}$ is equal to the cosmic baryon fraction
[…]
Using this condition, we solve for the baryon density amplitude $\rho_{0}$ to get
$\rho_{0}=\frac{\int_{0}^{r_{\rm b}}r^{2}\rho_{\rm NFW}(r)}{\int_{0}^{r_{\rm b}}r^{2}\left(\frac{x}{x_{\rm c}}\right)^{-\gamma}\left(1+\left(\frac{x}{x_{\rm c}}\right)^{1/\alpha}\right)^{-(\beta-\gamma)\alpha}}\;.$ (21)
This normalization depends on the baryon radius $r_{\rm b}$ where we expect the mass fraction in baryons to be $f_{\rm b}$. In other words, $r_{\rm b}$ represents the lower-bound of where the baryons are expected to trace the dark matter. We choose $r_{\rm b}=10\times r_{200}\approx 6.5\;\mathrm{Mpc}/h$ in comoving units. We note $r_{200}$ is the virial radius of the halo described in Eq. 14. The choice of $r_{\rm b}$ does impact $\rho_{0}$ if it is not chosen to be sufficiently large, for this reason we show it’s impact in Figure 6 and discuss this choice further in Appendix A. We note that this modelling choice does not explicitly preserve the mass of the halo (among other desirable properties) like the model framework of Bolliet et al. (2023), but it was shown by Cromer et al. (2022) that the inferred halo mass in galaxy cluster lensing simulations with baryons was unbiased with different choices of $r_{\rm b}$. In future work we anticipate modelling the entire profile using a halo model with galaxy-halo connection similar to works that model the imprint of the kSZ effect on the CMB power spectrum as a bispectrum (Smith et al., 2018; Münchmeyer et al., 2019; Cayuso et al., 2023; Roy et al., 2023; Bolliet et al., 2023). When doing a full halo model treatment, the choice of truncation radius is crucial and must be done rigorously so as to accurately capture the amount of gas contained in each halo, but for our work we do not truncate the 1-halo term as we are modelling a single effective halo which smoothly transitions into the 2-halo term through summation. We do not truncate the 1-halo term since it’s density amplitude will become negligible in comparison to the 2-halo density at sufficiently large radii $\sim 5\;\mathrm{Mpc}/h$ in comoving units.
3.2 The Two Halo Term
The kSZ effect probes the gas on the outskirts of halos thus we have to model the contribution of gas coming from neighbouring halos. Similarly, the GGL signal probes the excess surface mass density around galaxies across a wide range of scales from the virial radius of the central halo to scales hundreds of times the virial radius of the central halo. We use an analytical halo model to calculate the 2-halo term of the density profiles. The contribution to the total matter density profile from neighbouring halos is given by
$\rho^{\rm 2h}_{\rm m}(r)=\bar{\rho}_{\rm m}\xi^{\rm 2h}_{\rm hm}(r)A_{\rm 2h}\;,$ (22)
where $A_{\rm 2h}$ is an additional nuisance parameter to account for any potential systematic modelling errors in the 2-halo term, $\xi^{\rm 2h}_{\rm hm}(r)$ is the 2-halo term of the halo-matter correlation function, related to the matter correlation function by linear halo bias $\xi^{\rm 2h}_{\rm hm}=b(M)\xi^{\rm 2h}_{\rm mm}$, and $\bar{\rho}_{\rm m}$ is the comoving average cosmic matter density. We expect $A_{\rm 2h}$ to be less than unity to match the low amplitude of the 2-halo GGL signal Leauthaud et al. (2017); Amon et al. (2023) compared to predictions from clustering in Planck $\Lambda$CDM. As can be seen in Figure 2, the fit to only the kSZ effect measurements prefers much larger values of $A_{\rm 2h}$ than unity which is not physically justifiable. We choose our upper prior bound for $A_{\rm 2h}$ to minimize the prior volume effect (e.g. Ried Guachalla et al., 2025b) and reduce potential bias on the physical parameters of interest. We note that $M$ denotes the mass of each halo. The correlation function is related to the power spectrum by the inverse Fourier transform
$\xi_{\rm mm}^{\rm 2h}(r)=\int\frac{k^{2}dk}{2\pi^{2}}P^{\rm 2h}_{\rm mm}(k)j_{0}(kr)\;.$ (23)
The 2-halo term of the matter power spectrum $P^{\rm 2h}_{\rm mm}(k)$ is given by
[…]
where $P^{\rm lin}(k)$ is the linear matter power spectrum computed using the core cosmology library (pyccl) python package Chisari et al. (2019b), $dn/dM$ is the halo mass function Tinker et al. (2008), $b(M)$ is the linear halo bias Tinker et al. (2010), and $\hat{\rho}_{\rm NFW}(k)$ is the Fourier transform of the NFW profile. In this model we integrate over halo mass from $10^{10}M_{\odot}$ and $10^{15}M_{\odot}$. This completes the calculation of the 2-halo term of the total matter density profile. To compute the 2-halo contribution to the dark matter and baryon profiles we simply have
$\rho^{\rm 2h}_{\rm dm}(r)$ $=f_{\rm dm}\;\rho^{\rm 2h}_{\rm m}(r)$ (25)
$\rho^{\rm 2h}_{\rm b}(r)$ $=f_{\rm b}\;\rho^{\rm 2h}_{\rm m}(r)\;.$ (26)
This completes the calculation of the 2-halo term of the density profiles of matter around galaxies. We note that we assume the baryons trace the dark matter on large scales and that the gas is bound to the dark matter halos. For completeness, we write the complete density profiles for baryons and dark matter
$\rho_{\rm dm}(r)$ $=f_{\rm dm}\left(\rho_{\rm NFW}(r)+\bar{\rho}_{\rm m}\xi^{\rm 2h}_{\rm hm}(r)A_{\rm 2h}\right)$ (27)
$\rho_{\rm b}(r)$ $=f_{\rm b}\left(\rho_{\rm GNFW}(r)+\bar{\rho}_{\rm m}\xi^{\rm 2h}_{\rm hm}(r)A_{\rm 2h}\right)\;.$ (28)
In summary, the dark matter is described by an NFW profile and the baryons are described by a GNFW profile out to a radius of $r_{\rm b}=10\times r_{\rm 200}\approx 6.5\;\mathrm{Mpc}/h$ in comoving units. Beyond this radius the baryons trace the dark matter with cosmic mass fractions.
3.3 Modelling the kSZ Effect
For the galaxies in our lens sample which have a redshift range of $0.4\lesssim z\lesssim 0.7$, the optical depth is at the sub-percent level so $e^{-\tau}\approx 1$, which allows us to write
$\frac{\delta T_{\rm kSZ}}{T_{\rm CMB}}\approx\tau_{\rm gal}(\theta)\left(\frac{v_{\rm rms}^{\rm true}}{c}\right)\;.$ (29)
From cosmology and linear theory we know the RMS velocity of our lens sample is $v_{\rm rms}^{\rm true}\approx 1.06\times 10^{-3}c$ (Schaan et al., 2021; Amodeo et al., 2021), and $\tau_{\rm gal}$ is the optical depth to electrons of the galaxy group which can be written as
$\tau_{\rm gal}(\theta)=\sigma_{T}\int_{\rm LoS}\frac{d\chi}{1+z}n_{\rm e}\left(\sqrt{\left(\frac{\chi}{1+z}\right)^{2}+d_{\rm A}(z)^{2}|\theta^{2}|}\right)\;.$ (30)
Here we denote $d_{\rm A}(z)$ as the angular diameter distance of the galaxy group. With the kSZ effect we can probe extended gas in the outskirts of the dark matter halo. These measurements probe the electron density in the range of $\sim$1-4 virial radii and were first modelled by Amodeo et al. (2021). The minimum scale $\theta_{\rm min}\approx 1^{\prime}$ is dictated by the resolution of the stacked CMB maps while the maximum scale $\theta_{\rm max}\approx 6^{\prime}$ is chosen to minimize noise contamination from CMB primary anisotropies. At these scales the contribution to baryons from stars is negligible, therefore we say that $\rho_{\rm gas}(r)\approx\rho_{\rm b}(r)$. Furthermore, the ICM and CGM are broadly understood to be mostly ionized (McQuinn, 2016; Tumlinson et al., 2017), for this reason we assume that the gas is fully ionized with primordial abundances at these scales so that
$n_{\rm e}=\frac{X_{\rm H}+1}{2}\frac{\rho_{\rm gas}}{m_{\rm amu}}\;.$ (31)
We denote $X_{H}$ as the primordial mass fraction of hydrogen and $m_{\rm amu}$ to be an atomic mass unit. Combining all of this together we can say that given $\rho_{\rm gas}(r)$ we can model the kSZ temperature fluctuations to the CMB as
$\frac{\delta T_{\rm kSZ}(\theta)}{T_{\rm CMB}}=\left(\frac{v_{\rm rms}^{\rm true}}{c}\right)\frac{\sigma_{T}}{m_{\rm amu}}\frac{X_{H}+1}{2}\\ \times\underbrace{\int_{\rm LoS}\frac{d\chi}{1+z}\rho_{\rm gas}\left(\sqrt{\left(\frac{\chi}{1+z}\right)^{2}+d_{\rm A}(z)^{2}|\theta|^{2}}\right)}_{\rho^{\rm 2D}_{\rm gas}(\theta)}\;.$ (32)
The integral is a projection of the gas density profile along the line-of-sight. The temperature fluctuation maps $\delta T$ that are CAP filtered are observed through the ACT instrument beam, therefore we must convolve the projected density field with the real space beam profile $B(\theta)$ to make a direct comparison to the data. To perform this numerical convolution quickly and accurately, we use the Fast-Hankel Transform method described in Moser et al. (2023). The end result becomes
$\delta T_{\rm kSZ}(\theta)=T_{\rm CMB}\left(\frac{v_{\rm rms}^{\rm true}}{c}\right)\frac{\sigma_{T}}{m_{\rm amu}}\frac{X_{H}+1}{2}\left(\rho^{\rm 2D}_{\rm gas}(\theta)*B(\theta)\right)\;.$ (33)
With this expression for the kSZ effect temperature fluctuations of a model average galaxy of our sample, we can then do the same CAP filtering procedure described in Section 2.2 to compare to the measurements. This step is important for comparing the model to observations because there is some loss in signal caused by the CAP filter.
3.4 Modelling Galaxy-Galaxy Lensing
In GGL the measurement is of the excess surface mass density $\Delta\Sigma(R)$ which is fully described by the matter density profile $\rho_{\rm m}(r)$ previously described in Sections 3.1 and 3.2. The excess surface mass density $\Delta\Sigma(R)$ is a function of the comoving surface mass density $\Sigma(R)$ which we define as
[…]
where $\chi$ is the comoving line-of-sight distance to a redshift $z$ and $R$ refers to the projected comoving radius in the plane of the sky. [^4] Throughout this work $R$ refers to projected radii and $r$ refers to non-projected radii. The excess surface mass density $\Delta\Sigma$ is then defined as
$\Delta\Sigma(R)\equiv\overline{\Sigma}(\leq R)-\Sigma(R)\;,$ (35)
where the first term corresponds to the average comoving surface mass density enclosed within $R$ and is given by
[…]
We compute the total matter excess surface mass density using the FFTLog algorithm implemented in pyccl Chisari et al. (2019b). With this result we have a complete consistent framework for modelling the both the kSZ effect and GGL measurements.
4 Results: joint model constraints
Figure 2: 68% and 95% constraints on the halo model parameters (halo mass $M_{\rm halo}$, core scale fraction $x_{\rm c}$, outer power law slope $\beta$, and 2-halo amplitude $A_{\rm 2h}$) for the analysis of the BOSS kSZ temperature profile measurements (kSZ, red), GGL excess surface mass density profile measurements (GGL, blue) and joint observations (kSZ+GGL, purple). Jointly fitting both the kSZ and GGL measurements allows for tighter constraints on dark matter and gas parameters compared to fits using only one of the datasets, effectively "disentangling" the dark matter from the gas. The GGL measurements tightly constrain the mass of the sample, and probe the matter density out to much larger scales than the kSZ measurements, which effectively pins the amplitude of the 2-halo term $A_{\rm 2h}$ alleviating degeneracies in the other free parameters.
Figure 3: The baryon density profile (left), kSZ temperature profile at 150 GHz (top center) and 98 GHz (bottom center), and the GGL excess surface mass density profile (top right). We additionally show the fractional difference in $\Delta\Sigma$ relative to a model without any baryons (bottom right). The black points correspond to observations, the red corresponds to our fiducial model being fit to only the kSZ measurements, the blue corresponds to our fiducial model being fit to only the GGL measurements, and the purple corresponds to our joint fits to both the kSZ and GGL data. Dashed curves refer to the median (50th percentile) model of the posterior distribution and the bands denote the ±$\pm 1\sigma$ (32nd-68th percentiles) of the distribution of model curves. Lastly, solid curves denote the best-fit to the model. We show that the profile of the baryon density is tightly constrained by the joint fit relative to our kSZ or GGL only fits and qualitatively different compared to the data limited fits. We show that the joint fit of kSZ with GGL tightly constrains the baryon suppression in $\Delta\Sigma$ at small scales.
Here, we present a joint kSZ+GGL analysis of CMASS galaxies in the BOSS survey to disentangle the baryons and dark matter in the outskirts of the galactic halo. In our parameter fitting we assume a Gaussian likelihood for both measurements. The joint likelihood can be written as
$\ln\mathcal{L}\left[\vec{d}_{\rm kSZ},\vec{d}_{\rm GGL}\big{|}\vec{m}(\theta)\right]=-\frac{1}{2}\left(\vec{d}_{\rm kSZ}-\vec{m}(\theta)\right)^{T}\mathbf{C}^{-1}_{\rm kSZ}\left(\vec{d}_{\rm kSZ}-\vec{m}(\theta)\right)\\ -\frac{1}{2}\left(\vec{d}_{\rm GGL}-\vec{m}(\theta)\right)^{T}\mathbf{C}^{-1}_{\rm GGL}\left(\vec{d}_{\rm GGL}-\vec{m}(\theta)\right)\;.$ (37)
We refer to $\vec{d}_{\rm kSZ}$, $\vec{d}_{\rm GGL}$ as the measurements for the kSZ effect and GGL, and $\mathbf{C}_{\rm kSZ}$, $\mathbf{C}_{\rm GGL}$ as their respective covariance matrices, as described in Sections 2.1 and 2.2. One might expect that there would be non-negligible covariance between the two measurements because the kSZ probes the optical depth around galaxies while GGL probes the matter overdensity field around galaxies. We suspect that the optical depth could be correlated to the matter overdensity field and thus there could be non-negligible covariance for regions of the sky with significant overlap. In this analysis we assume the covariance between the kSZ and GGL measurements to be negligible due to a lack of overlap between the measurements on the sky (see Figure 4 of Schaan et al. (2021) and Figure 1 of Amon et al. (2023) for survey footprints of the kSZ and GGL measurements respectively).
We can write an expression for the posterior $\mathcal{P}$ as
$\ln\mathcal{P}(\theta\big{|}\vec{d}_{\rm kSZ},\vec{d}_{\rm GGL})=\ln\left[\mathcal{L}\left(\vec{d}_{\rm kSZ},\vec{d}_{\rm GGL}\big{|}\vec{m}(\theta)\right)\mathrm{Pr}(\theta)\right]\;.$ (38)
We refer to $\vec{m}(\theta)$ as our model function for both $T_{\rm kSZ}$ and $\Delta\Sigma$ given our model parameters $\theta$. We define $\mathrm{Pr}(\theta)$ to be the priors on our model parameters $\theta$ which can be found in Table 1. We sample the posterior using an affine-invariant ensemble Markov Chain Monte-Carlo (MCMC) algorithm implemented in the package emcee Foreman-Mackey et al. (2013). To ensure convergence we ran the chains sufficiently long such that the Gelman-Rubin convergence parameter $\mathcal{R}$ was less than 1.01 for all model parameters Gelman & Rubin (1992).
4.1 The complementarity of kSZ and GGL
We use the model that describes the two likelihoods to jointly fit for the mass of the dark matter halo $M_{\rm halo}$, the core scale fraction in the baryon gas profile $x_{\rm c}$, the outer power law slope for the baryon gas profile $\beta$, and the overall amplitude of the 2-halo term $A_{\rm 2h}$, which we include as a nuisance parameter. Figure 1 shows that the GGL measurements will be most sensitive to $M_{\rm halo}$ and $A_{\rm 2h}$ while the kSZ effect measurements will be most sensitive to the baryon gas parameters $x_{\rm c}$ and $\beta$. We also see that nearly all baryon gas parameters are either degenerate with $\beta$ or have negligible effects when varied. To alleviate some degeneracy we fix the inner power law slope $\gamma=0.2$, the intermediate power law slope $\alpha=1$.
Figure 2 shows the posteriors for our model parameters obtained from fitting to the stacked kSZ profile data (red). The best-fit (minimum $\chi^{2}$) for this model is $\chi^{2}_{\rm min}/\mathrm{d.o.f}=21.8/14$ (PTE $=0.08$). As one would expect, the halo mass and amplitude of the 2-halo term are largely unconstrained when the model only has access to the kSZ profile. [^5] Note that we tested the impact of using a substantially flexible prior on the amplitude of the 2-halo term, $A_{\rm 2h}=\mathcal{U}(0.0,20.0)$. In this case, we do constrain the parameter, $A_{\rm 2h}=3.2^{+1.4}_{-2.4}$. For our main analysis, we keep the physically-motivated prior to avoid exploring large parts of unphysical parameter space. Complementary are the blue contours, which show the posterior from the GGL data. The GGL measurements offer a lot of constraining power on $M_{\rm halo}$ and $A_{\rm 2h}$ but don’t have much constraining power for $\beta$ or $x_{\rm c}$. This fit to the data gave a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=2.5/7$ (PTE $=0.92$) implying that the GGL measurements were over-fit by our model. Furthermore, the GGL measurements are a probe of the total matter distribution which is comprised of mostly dark matter making it difficult to infer substantial information of the baryon density profile.
The joint fit to the data is shown in purple. This fit yielded a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=24.8/25$ (PTE $=0.47$) implying a good fit to the data. In these purple contours we can see the overlapping regions of unconstrained red and blue contours leading to relatively tight constraints on all parameters except $x_{\rm c}$. If we fix $x_{\rm c}=0.5$, then we can see that $M_{\rm halo}$, $\beta$, and $A_{\rm 2h}$ are all tightly constrained (denoted by dashed black line in Figure 2) with little degeneracy. By jointly fitting to both kSZ and GGL data we have ‘disentangled’ baryons from dark matter.
Figure 3 shows the best-fit models from the kSZ, GGL, and kSZ+GGL data for the baryon density profile $\rho_{\rm b}(r)$ as a function of radius (left), the kSZ profile (middle) and the GGL (right). The kSZ measurements constrain the density on small scales near the transition from the 1-halo to 2-halo term and the GGL measurements are good at constraining $\rho_{\rm b}(r)$ at larger scales in the 2-halo term. When combined we see a tight constraint on the density profile across all scales, within our model where $\alpha=1$ and $\gamma=0.2$ are fixed. We explored allowing $\alpha$ to be a free parameter within the model and found that the density profile is still well constrained across all scales despite $\alpha$ being unconstrained. Having tight constraints on the baryon density profile across such a wide range of scales enables more direct comparisons to simulations.
In the bottom right panel of Figure 3 we show the fractional difference in the GGL signal due to baryons $\delta(\Delta\Sigma)/\Delta\Sigma=(\Delta\Sigma-\Delta\Sigma_{\rm nb})/\Delta\Sigma_{\rm nb}$, where $\Delta\Sigma_{\rm nb}$ refers to the same fiducial model except we exclude baryons ($f_{\rm b}=0\;\rightarrow\;f_{\rm cdm}=1$). When the model only has access to the GGL measurements, the suppression in $\Delta\Sigma$ at small scales is unconstrained but when we include the kSZ profile measurements into the fit we can clearly see that suppression in $\Delta\Sigma$ at small scales is well constrained.
Figure 4: Constraints on the halo mass from GGL where baryons are neglected (dark blue, $f_{\rm cdm}=1$), where the baryon contribution and its uncertainty is modelled (blue, $f_{\rm cdm}=1-f_{\rm b}$) and where kSZ is jointly fit with the GGL to constrain the baryon parameters (purple). The prior range for the halo mass adopted by (Bigwood et al., 2024) to encompass the uncertainty on the stellar masses and the SHMR spans beyond the extent of the plot.
4.2 Halo Mass
As a consequence of the suppression of $\Delta\Sigma$ caused by baryons, as seen in Figure 3 (lower right), neglecting to account for baryons can bias halo mass estimates $M_{\rm halo}$, compared to the true mass (e.g. Debackere et al., 2020; Cromer et al., 2022). This effect is demonstrated in Figure 4 where we show the posteriors of the halo mass for three analyses of the GGL data: fitting a model that neglects baryons ($f_{\rm b}=0\;\rightarrow\;f_{\rm cdm}=1$) to the GGL data, fitting our model which includes baryons to the same GGL data, and the same, where we jointly analyse GGL with kSZ profile measurements. We find that neglecting the impact of baryons lowers our estimate of the halo mass by $\sim 20\%$. Including the impact of baryons in the model but only using GGL data inflates our posterior, which is improved by the inclusion of the kSZ measurements. That is, a joint analysis of GGL and kSZ can account for the gas parameters, shrinking the width of the posterior back down, while the halo mass estimate undergoes a $\sim 20\%$ shift compared to the no-baryon model. We note that our joint kSZ+GGL constraints on $M_{\rm halo}$ are expected to be underestimated by $\lesssim 10\%$ compared to the mean halo mass of the sample $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$ likely due to the exclusion of miscentering in our model (Skibba et al., 2011; Hoshino et al., 2015).
A crucial step in using kSZ profile measurements as a benchmark for feedback models and simulations is an accurate constraint on the halo mass. The halo mass can change both the shape and the amplitude of the kSZ signal as seen in Figure 1. The ratio of kSZ signals modelled with halo masses of $10^{14}\;\rm M_{\odot}$ and $10^{13}\;\rm M_{\odot}$ is $\approx 2.1$ at $\theta=1^{\prime}$ and grows with increasing $\theta$, plateauing to $\approx 3.3$ at $\theta=6^{\prime}$. Assuming a certain halo mass estimated from stellar masses is potentially dangerous as stellar masses and SHMR can be biased (Yuan et al., 2021; Oyarzún et al., 2024). These kSZ profile measurements were jointly analysed with cosmic shear for the first time by Bigwood et al. (2024) to constrain the power spectrum and they found a best-fitting model with stronger baryonic feedback than hydrodynamical simulations predict. As discussed in Bigwood et al. (2024) (their Appendix B3), a major source of uncertainty in constraining the impact of baryonic feedback from kSZ and cosmic shear measurements, was the halo mass of the sample, which spans almost an order of magnitude: $M_{\rm halo}=0.8-30\times 10^{13}\;M_{\odot}$ (significantly larger than the range of masses plotted in Figure 4). For this reason, Bigwood et al. (2024) had to use a large prior for $M_{\rm halo}:\mathcal{U}[5\times 10^{12},7\times 10^{13}]$. Having such a large uncertainty on halo masses make it difficult to interpret the strength of baryonic feedback from the stacked kSZ profiles measured by Schaan et al. (2021); Hadzhiyska et al. (2024a) and future measurements to come. The methodology in this work can provide a more precise prior on the effective halo mass of the sample.
We compare our ($2\sigma$) posteriors on the halo mass $\log_{10}(M_{\rm halo,200m}/\rm M_{\odot})=13.44\pm 0.12$ to the halo mass found by McCarthy et al. (2024) where they jointly analysed the kSZ and GGL measurements with the FLAMINGO simulation suite (Schaye et al., 2023). In this work we’ve used a mass definition of $M_{\rm 200m}$, to match McCarthy et al. (2024) we transform to a mass definition of $M_{\rm 500c}$ using an NFW profile (Navarro et al., 1997) and the concentration mass relation of Duffy et al. (2008). Our ($2\sigma$) posteriors on the halo mass are $\log_{10}(M_{\rm halo,500c}/\rm M_{\odot})=13.22\pm 0.12$ and the ($2\sigma$) posteriors on the halo mass from McCarthy et al. (2024) are $\log_{10}(M_{\rm 500c}/\rm M_{\odot})=13.34\pm 0.04$. We find that our estimate of the halo mass is $\sim 30\%$ lower than what was found by McCarthy et al. (2024).
Moreover, getting an unbiased estimate of the halo mass is crucial for modelling the tSZ effect (Moser et al., 2021, 2023) and the matter power spectrum suppression due to baryons (Debackere et al., 2020). Unlike the kSZ effect which scales linearly with halo mass $T_{\rm kSZ}(\theta)\propto M_{\rm halo}$, the tSZ effect has a much stronger scaling with halo mass $T_{\rm tSZ}(\theta)\propto M_{\rm halo}^{5/3}$. For this reason, an estimated halo mass that is biased high by $30\%$ will bias the model to over predict the amplitude of $T_{\rm tSZ}(\theta)$ by at least $50\%$. This potential bias arising from incorrect halo mass estimates could help alleviate discrepancies between the observed and predicted tSZ profiles in Moser et al. (2022) and tSZ clustering signals described in McCarthy et al. (2024). Similarly, a bias in halo mass estimates of $\sim 30\%$ could lead to an underestimation of the suppression of the matter power spectrum by $\sim 4\%$ at $k=1\;h/\mathrm{Mpc}$ (Debackere et al., 2020). By virtue of having an analytic model that can consistently model both the kSZ profile measurements and the GGL measurements we can completely side step any assumptions made to estimate the halo mass of the sample.
In future work we will improve our model by including satellites and accounting for miscentering which will yield an entirely unbiased estimate of the halo mass.
4.3 Baryonic Feedback
The kSZ effect probes the distribution of gas in the outskirts of halos. Only recently have direct measurements of the stacked kSZ effect profile had high enough SNR to be a powerful probe of the gas (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a). The ACT-BOSS kSZ measurements used in this analysis from Schaan et al. (2021) were first analysed with their corresponding tSZ measurements to study the thermodynamic properties of the halo by Amodeo et al. (2021) and have recently been jointly analysed with cosmic shear for the first time by Bigwood et al. (2024) to constrain the matter power spectrum. The joint analysis of Bigwood et al. (2024) found a best-fitting model that preferred a more extreme baryonic feedback scenario than hydrodynamical simulations predict. These same kSZ and GGL measurements were then analysed using the FLAMINGO simulations directly by McCarthy et al. (2024). Their analysis found that the kSZ measurements required more extreme baryonic feedback than the FLAMINGO simulations predicted by $7\sigma$. With this context in mind, we compare our joint kSZ+GGL constraints to the GNFW gas profile scaling relations derived from hydrodynamical simulations in Battaglia (2016).
It can be seen in Figure 1 that the outer power law slope $\beta$ is the most important parameter in our model for describing the distribution of baryons in the halo outskirts. Since we are not modelling the entire distribution of galaxies with an HOD, we are measuring an effective value of $\beta$ which inherently assumes that all halos in our sample have the same mass. This effective $\beta$ could be different than an HOD-based model would find and we leave this as a future improvement to the model. In Figure 5 we compare our constraints on $\beta$ to scaling relations $\beta(M,z)$ derived from the cosmological hydrodynamic simulations of Battaglia (2016) (corrected for difference in GNFW profile definition, see Appendix B). The $\beta(M,z)$ relation was derived from gas profiles of resolved halos with masses of $M_{500}=10^{14}-10^{15}\;\rm M_{\odot}$, for this reason we acknowledge that our comparisons are to an extrapolation of the $\beta(M,z)$ relation of Battaglia (2016).
We show our ±$\pm 1\sigma$ estimates on $\beta$ and $M_{\rm halo}$ compared to 2 models from Battaglia (2016). The blue curve in Figure 5 is the Adiabatic model which corresponds to a simulation with only gravitational heating, the red curve is the AGN model which includes radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback (more extreme feedback). A lower value of $\beta$ implies that the baryons are more spread out to larger radii implying stronger baryonic feedback. We find that our posteriors on halo mass and $\beta$ are consistent with the assumptions and findings of Amodeo et al. (2021) where they found $\beta=2.6^{+1.0}_{-0.6}$ (at fixed mean halo mass $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$: light blue in Figure 5 with necessary correction discussed in Appendix B) when fitting a GNFW gas profile to the same kSZ measurements. [^6] We note that the modelling done in Amodeo et al. (2021) includes a mass-weighted average over the halo masses in their sample, for comparison we plot their constraints on $\beta$ at the mean halo mass in their sample in Figure 5. We find our constraints on $\beta$ are tighter than Amodeo et al. (2021) by a factor of 2 when fitting our model to the same data (kSZ Only fit in Table 1). This improvement in constraining $\beta$ is due to the way we normalize the baryon (gas) density profile (see Eq. 21) which effectively increases the importance of $\beta$ in the model. When we include GGL in the kSZ+GGL joint analysis our constraints on $\beta$ tighten further by 50% yielding a total factor of 3 improvement compared to Amodeo et al. (2021) while also giving a direct estimate on the halo mass $M_{\rm halo}$.
When fitting our fiducial model to kSZ+GGL measurements, we find that our constraints on $\beta$ are lower than the AGN cosmological hydrodynamic simulation feedback model of Battaglia (2016). We note that our constraints on $\beta$ do change with respect to our choice of baryon radius $r_{\rm b}$ described in Appendix A. In Figure 5, we also show our fiducial choice of $r_{\rm b}=10\times r_{200}$ compared to an extreme choice of $r_{\rm b}=20\times r_{200}$. The baryon radius $r_{\rm b}$ represents the radius at which the mass fraction of baryons to matter is equal to the cosmic baryon fraction $f_{\rm b}$ and it is crucial for normalizing the amplitude of the baryon density profile $\rho_{0}$. As $r_{\rm b}$ increases $\beta$ does too, but if the radius is too large the applicability of the NFW profile used to describe the dark matter and the GNFW profile used to describe the baryons is no longer appropriate. We show the impact of our choice of $r_{\rm b}$ on kSZ+GGL posteriors in Figure 6, but in all cases our joint kSZ+GGL analysis implies that baryonic feedback is stronger than the cosmological hydrodynamic simulations of Battaglia (2016) with AGN feedback predict.
In Figure 3 we compare our results to hydrodynamical simulations where available. In the rightmost panel we show the suppression of $\Delta\Sigma$ in prominent hydrodynamical simulations: Illustris (Vogelsberger et al., 2014) (dark red, dot-dashed) and IllustrisTNG-300 (Pillepich et al., 2018) (dark red, dotted). These curves are from Lange et al. (2019) where they show the impact of baryonic feedback on the GGL signal around CMASS-like halos with a stellar mass range of ($\log_{10}(M_{*}/\rm M_{\odot})=10.5-12$) in Illustris and IllustrisTNG. They select halos from snapshots matching the mean redshift of the CMASS sample $\langle z\rangle=0.55$ and compute the suppression in $\Delta\Sigma$ from baryons for 4 separate stellar mass bins. For a given simulation (Illustris or IllustrisTNG), the difference in suppression within $R\lesssim 1\;\mathrm{Mpc}/h$ (in comoving units) is within a few percent. In Figure 3 we plot the average of the $11\leq\log_{10}(M_{*}/\rm M_{\odot})\leq 11.5$ bin and the $11.5\leq\log_{10}(M_{*}/\rm M_{\odot})\leq 12.0$ to most closely represent the expected suppression of our sample.
[…]
Our last point of comparison to simulations is directly to the baryon (gas) density profile shown in the leftmost panel of Figure 3. Alongside our fits, we show curves from Amodeo et al. (2021) for the AGN feedback model of Battaglia (2016) (green, dash-dot) also depicted in Figure 5 (red, solid) and the IllustrisTNG simulations (dark red, dotted). To produce the density profile for the AGN feedback simulation, Amodeo et al. (2021) used the best-fit relationships from Battaglia (2016) to compute a mass-weighted density profile with weights chosen to match the CMASS halo mass distribution. They chose the redshift to be the mean redshift of the CMASS sample $\langle z\rangle=0.55$. For the IllustrisTNG density profile, they modelled the profile by selecting “red” galaxies with colours sdss_g - sdss_r $\geq$ 0.6 (Nelson et al., 2018) and weighting each galaxy by stellar mass to match the corresponding stellar mass distribution of the CMASS sample. We see that the two simulations are similar in shape and amplitude and appear to be comparable to our fits to the measurements. A deeper exploration of modelling choices on sample selection in the simulations was done by Moser et al. (2021), and they found for the IllustrisTNG simulations with the same mock selection as done in Amodeo et al. (2021) that $\beta=3.20\pm 0.11$ ($\beta=2.8\pm 0.11$ when corrected for difference in profile definition, see Appendix B). This value of $\beta$ sits between the AGN model of Battaglia (2016) and our kSZ+GGL fits with $r_{\rm b}=10\times r_{200}$. It would be informative to have similar GNFW density profile parameter fits found in Battaglia (2016) from different hydrodynamical simulations to get a better understanding of how our measurements compare to other feedback schemes.
We note that our model neglects satellites, for this reason we are neglecting the potential systematic of ’miscentering’. We expect that not accounting for this effect causes an underestimation of the kSZ amplitude in simulations by up to $\sim 20\%$ for CMASS-like galaxies ($12.111.5<\log M_{*}<$ 12.1) as shown in Figure 6 of Hadzhiyska et al. (2023), at the current level of statistical power, this effect would not significantly change the conclusions of this work. We note that this effect is dependent on the satellite fraction of the sample which means that miscentering is more significant for lower mass galaxies ($11.510.9<\log M_{*}<$ 11.5) and less significant for more massive galaxies ($12.712.1<\log M_{*}<$ 12.7). Considering Figure 1, we can see that ignoring miscentering from satellites in our model would most likely mean our values for $\beta$ are overestimated by $\sim 0.2-0.4$ which would not significantly impact the conclusions from this study. There is of course, a possibility that miscentering could alter the shape of the kSZ profile further altering the value of $\beta$ changing our results. Including this effect in the model can become computationally expensive, but it has been modelled for observations of the tSZ effect (Popik et al., 2025). We leave this improvement to our model of kSZ+GGL as future work.
Figure 5: Scaling relations derived from Battaglia (2016) hydrodynamical simulations for the outer power law slope, $\beta$, of the baryon density profile in our model as a function of halo mass, $M_{\rm halo}$. The blue line corresponds to simulations with only gravitational heating and the red line corresponds to simulations with additional radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback. Lower values of the outer power law slope $\beta$ indicate the baryons are more diffusely spread out in the density profile to the outermost regions of the halo. This suggests that the data prefers stronger baryonic feedback relative to the hydrodynamical simulations of Battaglia (2016). In our model we use a fiducial value for the baryon radius $r_{\rm b}=10\times r_{200}$ (purple), but we include an additional data point with $r_{\rm b}=20\times r_{200}$ to show the range of values $\beta$ could be with this model depending on the choice of $r_{\rm b}$.
5 Conclusions
In this work we present an analytic framework to model GGL measurements jointly with stacked kSZ profile measurements to disentangle the halo. We jointly analyse combined GGL measurements from DESY3 and KiDS-1000 to probe the matter density around BOSS CMASS galaxies Amon et al. (2023), while our stacked kSZ profile measurements from ACT DR5 and Planck maps probe the gas density around the same galaxies Schaan et al. (2021). Jointly modelling these two signals offers great potential for disentangling baryons from dark matter in the outskirts of dark matter halos at the expense of requiring a complex model to consistently describe both signals which (to our knowledge) has never been done before. We model the signals using an analytical halo model to fit for both the halo mass and the gas density profile parameters simultaneously. Our most important findings from our analysis are:
  • I.
[…]
  • II.
    Given our modelling choices and assumptions, we find our model is consistent with enhanced baryonic feedback compared to the hydrodynamical simulations from Battaglia (2016). This comes from comparing our kSZ+GGL constraints of the outer power law slope $\beta$ on the baryon density profile to relationships $\beta(M,z)$ derived from Battaglia (2016) hydrodynamical simulations, as seen in Figure 5. Furthermore, we find that our kSZ+GGL constraints on the suppression in $\Delta\Sigma$ due to baryons is better represented by the Illustris simulation than the IllustrisTNG simulation (see Figure 3).
Our model neglects the modelling of satellites which can have a significant impact on the kSZ profile through miscentering and could bias our inferred effective value of $\beta$. The next step forward in improving the model would be to account for the galaxy-halo connection using an HOD framework similar to the work done by Popik et al. (2025) for the tSZ effect. This improvement would enable us to account for the presence of satellites and it enables a self-consistent halo model treatment of the 1-halo and 2-halo terms for an accurate model of the matter spectrum. This would allow the kSZ measurements to be used a cosmological analysis of galaxy-galaxy lensing and clustering to model the impact of baryon feedback on smaller scale measurements.
Going forward, there are many avenues to build upon this analysis. First, new measurements of the stacked kSZ profile from ACT-DR6 and DESI-Y1 data using photometric (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2024; Hadzhiyska et al., 2024b) and spectroscopic (Ried Guachalla et al., 2025a) redshifts have significantly higher SNR than the measurements from Schaan et al. (2021) (analysed in this work) with the ability to measure the signal in a more narrow range of redshift and mass. In the near future, Simons Observatory (SO) will bring better CMB measurements to accompany the DESI spectroscopic survey and yield better kSZ profile measurements. At the same time, upcoming WL surveys such as Euclid, LSST, and Roman will provide substantially more powerful WL measurements. Additionally, when combined, the kSZ effect and the tSZ effect can probe the thermodynamic profiles around galaxies (Amodeo et al., 2021), when combined with additional information from GGL one can have unprecedented constraining power for testing baryonic feedback models. Together, these measurements will enable a better understanding of baryonic feedback, halo masses, and cosmology.
Acknowledgements
We would like to thank Nick Kokron, Gabriela Sato-Polito, Dhayaa Anbajagane, Henry Schreiner, Adrian Bayer, and the members of the Amon research group for insightful discussions. We are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is a consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Research Computing. JS acknowledges and is thankful for the support from the Fannie & John Hertz Foundation. JD acknowledges support from a Royal Society Wolfson Visiting Fellowship and from the Kavli Institute for Cosmology Cambridge and the Institute of Astronomy, Cambridge.
Data Availability
The stacked kSZ profile measurements used in this work can be made available upon request and are from the original work of Schaan et al. (2021). The GGL measurements used in this analysis are publicly available on this github repository and were originally made by Amon et al. (2023). [^7] https://github.com/aamon/Galaxy-galaxy-lensing-KiDS-DES-HSC Our modelling code glasz is built upon the pyccl Chisari et al. (2019b) and Mop-c-GT Amodeo et al. (2021) packages and is publicly available at the glasz homepage. [^8] https://github.com/James11222/glasz We request that any uses of the glasz package cite the DOI listed on the package homepage.
Appendix A Impact of Baryon Radius
We define the baryon radius $r_{\rm b}$ to be the radius where we expect the mass fraction in baryons to be the cosmic fraction $f_{\rm b}$. In our analysis, we chose $r_{\rm b}=10\times r_{200}$, but the choice of this radius does shift the posteriors of our model parameters. One can see the exact behaviour in Figure 6 where we show what posteriors look like for four choices of $r_{\rm b}$ with increasing size. The most notable effect of increasing $r_{\rm b}$ is that the outer power law slope $\beta$ of the baryon density profile also increases. We note that this shift in $\beta$ does begin to saturate at sufficiently large $r_{\rm b}\gtrsim 10\times r_{200}$. In all cases, our posteriors of $\beta$ still suggest values less than to be expected by the cosmological hydrodynamic simulations with AGN feedback used in Battaglia (2016) as mentioned in Section 4.
If $r_{\rm b}$ is chosen to be too small, then we see a bias in $A_{\rm 2h}$. This bias occurs because we don’t enforce normalization of the baryon density profile amplitude $\rho_{0}$ sufficiently far out causing $\beta$ to compensate creating an outer power law slope that is comparable to or shallower than an NFW profile. If the baryon profile is shallower than the dark matter profile then one would expect the baryon density 1-halo term to be non-physically large at large scales forcing $A_{\rm 2h}$ be biased low to compensate for this.
Figure 6: We show 4 different choices of $r_{\rm b}$ ranging from $r_{\rm b}=5\times r_{200}$ to $r_{\rm b}=20\times r_{200}$ to demonstrate the impact of the choice of $r_{\rm b}$ on our posterior for dark matter and baryon profile parameters. As $r_{\rm b}$ increases we note that the outer power law slope $\beta$ increases while the halo mass $M_{\rm halo}$ and 2-halo amplitude $A_{\rm 2h}$ remain relatively unchanged. We note that as $r_{\rm b}$ becomes sufficiently large $r_{\rm b}\gtrsim 10\times r_{200}$, it’s affect on $\beta$ saturates in our fits. If $r_{\rm b}$ is chosen to be too large, then the validity of the single halo radial profiles used in our model breaks down. If $r_{\rm b}$ is chosen to be too small (i.e $r_{\rm b}=5\times r_{200}$) then we add risk of assuming universal mass fraction prematurely which implies lower values of $\beta$ potentially implying stronger unphysical baryonic feedback. For these reasons, our fiducial model uses $r_{\rm b}=10\times r_{200}\approx 6.5\;\mathrm{Mpc}/h$ in comoving units.
Appendix B Differences from Amodeo Halo Model
In this section we unpack a few important differences from our model and pre-existing work from Amodeo et al. (2021) that directly precedes ours.
  • We use a slightly different GNFW form for the density profile than Eq. 16 of Amodeo et al. (2021). Our version follows the original form of Zhao (1996). As mentioned in Smith et al. (2018), the GNFW form used in (Battaglia, 2016; Amodeo et al., 2021) must be corrected to be consistent with the universal NFW profile proposed in Zhao (1996). The difference between the two profiles amounts to $\Delta\beta=2|\gamma|$ assuming fixed $\alpha=1$. More specifically, any $\beta$ which uses the GNFW profile of (Battaglia, 2016; Amodeo et al., 2021) will be larger by $\Delta\beta=0.4$ assuming a fixed $|\gamma|=0.2$ where the sign is chosen so that all radial terms in the profile are in the denominator. While correcting for this does bring our joint fits closer to the simulated models for $\beta(M,z)$ from Battaglia (2016), it does not substantially change the results of this work. In all comparisons of $\beta$ to other works (e.g. Battaglia, 2016; Amodeo et al., 2021; Moser et al., 2021) we account for this necessary correction.
  • Our framework removes $\rho_{0}$ as a free parameter from the model of the baryon density profile by forcing a normalization of the profile at $r_{\rm b}$. This normalization procedure was inspired by Cromer et al. (2022). Without this normalization the model prefers the amplitude of baryons to be an order of magnitude lower than required by cosmological abundance.
  • The Mop-c-GT code used in Amodeo et al. (2021) has dependencies on two cosmology codes HMF and colossus which have different assumptions of units. These different assumptions do not appear to be properly accounted for in the 2-halo term module of Mop-c-GT. For this reason we found that the 2-halo density curve used in Amodeo et al. (2021) was a factor of $1/h^{2}$ larger than it should have been. While $A_{\rm k2h}$ was a nuisance parameter and was marginalised over, this finding implies that their posteriors for $A_{\rm k2h}$ should be $1/h^{2}$ larger. Our implementation of this model in glasz uses pyccl for all halo model computations rather than combining multiple cosmological codes.

### v2

@@
 Abstract
-[…]
+We present the first joint analysis of the kinetic Sunyaev-Zeldovich (kSZ) effect with galaxy-galaxy lensing (GGL) for CMASS galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS).
+We show these complementary probes can disentangle baryons from dark matter in the outskirts of galactic halos by alleviating model degeneracies that are present when fitting to kSZ or GGL measurements alone.
+In our joint kSZ+GGL analysis we show that the baryon density profile is well constrained on scales from 0.3 to 50 Mpc/$h$.
+With our well constrained profile of the baryon density, we provide direct comparisons to simulations.
+For our model we find an outer slope of the baryon distribution that is shallower than predicted by some hydrodynamical simulations, consistent with enhanced baryonic feedback.
+We also show that not including baryons in a model for GGL can shift halo mass estimates by $\sim 20\%$ compared to a model that includes baryons and is jointly fit to kSZ+GGL measurements.
+Our modelling code galaxy-galaxy lensing and kSZ (glasz) is publicly available at https://github.com/James11222/glasz.
 keywords:
 gravitational lensing: weak – galaxies: haloes – cosmology: observations – (cosmology:) cosmic background radiation
-pubyear: 2015pagerange: Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy LensingB
+pubyear: 2026pagerange: Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy LensingC
 1 Introduction
@@
 One of the biggest challenges standing in the way of using such measurements to test the dark components of the standard model is the modelling of baryons.
-In particular, baryonic feedback suppresses the matter power spectrum on non-linear scales, and therefore the weak lensing amplitude, and must be modelled accurately to avoid biasing cosmological inference (e.g., van Daalen et al., 2011a; Chisari et al., 2019a; Bigwood et al., 2024).
+In particular, baryonic feedback suppresses the matter power spectrum on non-linear scales, and therefore the weak lensing amplitude, and must be modelled accurately to avoid biasing cosmological inference (e.g., van Daalen et al., 2011b; Chisari et al., 2019b; Bigwood et al., 2024).
 However, the complexity of baryonic feedback requires a large dynamic range in resolution and time that is difficult to simulate (e.g.
-Crain & van de Voort, 2023).
-This has led to the development of simplified ‘subgrid’ prescription models in hydrodynamical cosmological simulations which are used to mimic the effects of baryonic feedback on LSS (Steinborn et al., 2015; Chisari et al., 2019a; Schaye et al., 2023; Pakmor et al., 2023).
-Although the simulations are able to reproduce a range of observations, they give a wide spread of predictions for the amplitude and scale extent of the suppression of the matter power spectrum due to baryonic feedback (e.g., Chisari et al., 2019a; Villaescusa-Navarro et al., 2021; Sunseri et al., 2023).
+Crain and van de Voort, 2023).
+This has led to the development of simplified ‘subgrid’ prescription models in hydrodynamical cosmological simulations which are used to mimic the effects of baryonic feedback on LSS (Steinborn et al., 2015; Chisari et al., 2019b; Schaye et al., 2023; Pakmor et al., 2023).
+Although the simulations are able to reproduce a range of observations, they give a wide spread of predictions for the amplitude and scale extent of the suppression of the matter power spectrum due to baryonic feedback (e.g., Chisari et al., 2019b; Villaescusa-Navarro et al., 2021; Sunseri et al., 2023).
 Understanding and reducing this model uncertainty is the challenge.
@@
 Another approach is to jointly model the WL and clustering data with probes of the baryon content (Debackere et al., 2020; Tröster et al., 2022; Schneider et al., 2022; Bigwood et al., 2024).
-The kinetic Sunyaev-Zeldovich (kSZ) effect is a secondary anisotropy in the CMB that comes from the Doppler boost of CMB photons by inverse Compton scattering of free electrons in the Intracluster Medium (ICM) and Circumgalactic Medium (CGM) that have bulk motion along the line of sight relative to the CMB rest frame (Zeldovich & Sunyaev, 1969; Mroczkowski et al., 2019).
+The kinetic Sunyaev-Zeldovich (kSZ) effect is a secondary anisotropy in the CMB that comes from the Doppler boost of CMB photons by inverse Compton scattering of free electrons in the Intracluster Medium (ICM) and Circumgalactic Medium (CGM) that have bulk motion along the line of sight relative to the CMB rest frame (Zeldovich and Sunyaev, 1969; Mroczkowski et al., 2019).
 The kSZ effect is small relative to other CMB secondaries (e.g.
@@
 The dependency on the bulk line of sight velocity and the small amplitude relative to other secondaries requires methods such as a velocity weighted stack using external velocity measurements from galaxy surveys to measure the signal.
-Stacked measurements with a velocity reconstruction weighting have revealed that the electrons around galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS) are significantly more extended than the dark matter distribution (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a).
+Stacked CMB measurements from the Atacama Cosmology Telescope (ACT) survey with a velocity reconstruction weighting have revealed that the electrons around galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS) are significantly more extended than the dark matter distribution (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b).
 The kSZ effect is particularly useful for WL as it probes the density of electrons (and therefore baryons) in the outskirts of the halo ($\sim$1-4 virial radii) which lies in the range of scales that baryons need to be better understood to improve WL cosmological constraints.
@@
 By jointly analysing the kSZ effect with GGL measurements we can build a self-consistent framework for estimating the halo mass.
-Without GGL, analyses of the kSZ effect either rely on estimates of stellar mass (Maraston & Strömbäck, 2011; Maraston et al., 2013) which are then used to estimate the halo mass through stellar-to-halo mass relations (Miyatake et al., 2015; Kravtsov et al., 2018), or fits to galaxy clustering (Miyatake et al., 2015; Rodríguez-Torres et al., 2016), both of which have greater uncertainties than galaxy-galaxy lensing estimates.
+Without GGL, analyses of the kSZ effect either rely on estimates of stellar mass (Maraston and Strömbäck, 2011; Maraston et al., 2013) which are then used to estimate the halo mass through stellar-to-halo mass relations (Miyatake et al., 2015; Kravtsov et al., 2018), or fits to galaxy clustering (Miyatake et al., 2015; Rodríguez-Torres et al., 2016), both of which have greater uncertainties than galaxy-galaxy lensing estimates.
 Bigwood et al.
@@
 (2021) are consistent with feedback that is stronger than most hydrodynamical simulations predict (Bigwood et al., 2024; McCarthy et al., 2024).
-This underestimation may be severe enough that it may even bias WL cosmological constraints (Amon & Efstathiou, 2022; Preston et al., 2023).
+This underestimation may be severe enough that it may even bias WL cosmological constraints (Amon and Efstathiou, 2022; Preston et al., 2023).
 In this work we build a fast analytic modelling framework based on the halo model, complementary to previous approaches that rely on the baryonification method, or expensive hydrodynamical simulations.
-This framework can be used for cosmological analyses of GGL and clustering in the future.
+While out of the scope of this work, validation of our model with simulations will be necessary to use this model for cosmological analyses of GGL and clustering in the future.
 This work is structured as follows: Section 2 describes the kSZ and GGL measurements that we jointly analyse, Section 3 details the model that we construct, Section 4 demonstrates the model in practice and presents the joint fits for BOSS galaxies and Section 5 concludes the work.
@@
 In this work, we build a model to consistently describe the excess surface mass density of halos, as measured by GGL (Amon et al., 2023), with the gas distribution, probed by measurements of the kSZ effect (Schaan et al., 2021).
+These measurements require data from a combination of several surveys (BOSS, DES Y3, KiDS-1000, ACT DR5), for clarity we summarize the sky footprints and redshift distributions of these surveys in Figure 1.
 We describe these measurements in more detail below.
+Figure 1: Left: Sky footprints in equatorial coordinates of all surveys used in this work.
+Spectroscopic lenses are from the BOSS CMASS (purple) survey, photometric sources used for weak lensing come from DES Y3 (red) and KiDS-1000 (black), and the CMB backlight used for stacking the kSZ signal is from the ACT DR5 (orange) survey.
+We mask the full ACT footprint using the Planck $f_{\rm sky}=60\%$ galactic mask from Planck Collaboration et al.
+(2020a).
+We estimate that DES Y3, KiDS-1000, and ACT have an overlap area of 777 deg2, 409 deg2, and 5737 deg2 with BOSS respectively.
+Right: Redshift distributions for the galaxy surveys used in this analysis from Amon et al.
+(2023) (BOSS scaled by $10^{3}$).
+The BOSS CMASS survey is divided into 2 samples C1 (pink) and C2 (purple) by redshift.
+DES Y3 and KiDS-1000 survey bins are coloured to indicate which lens bin they correspond to, such that the sources are sufficiently behind the lenses.
 2.1 Galaxy-Galaxy Lensing
@@
 (2021).
-The KiDS-1000 weak lensing data were taken with the 2.6 m VLT Survey Telescope (VST) (Capaccioli & Schipani, 2011; Grado et al., 2012) and span 777 $\rm deg^{2}$ of the sky (Giblin et al., 2021) with a number density of 6.22 $\rm arcmin^{-2}$.
+The KiDS-1000 weak lensing data were taken with the 2.6 m VLT Survey Telescope (VST) (Capaccioli and Schipani, 2011; Grado et al., 2012) and span 777 $\rm deg^{2}$ of the sky (Giblin et al., 2021) with a number density of 6.22 $\rm arcmin^{-2}$.
+The spectroscopic lenses from BOSS CMASS are subdivided into 2 lens redshift bins: C1 corresponding to $0.43<z<0.54$ and C2 corresponding to $0.54<z<0.70$.
+Each lensing survey (DES Y3 and KiDS-1000) providing background source galaxy shapes is also divided into 4 tomographic redshift bins, of which, we only use tomographic redshift bins sufficiently behind the foreground lens redshift bins (a GGL measurement of lens bin C1 uses both tomographic bins [3,4] while C2 uses only the [4] tomographic bin).
+The redshift distributions $n(z)$ of the lenses and sources are shown in the top-right panel of Figure 1.
+The tomographic bins used in DES Y3 and KiDS-1000 are calibrated using a tool from machine-learning known as the self-organizing map (SOM) (Buchs et al., 2019; Wright et al., 2020) which requires spectroscopy to accurately estimate photometric redshifts in each tomographic bin.
 Given spectroscopically measured redshifts of lenses and photometric shape and redshift measurements of the sources, the quantity that is directly measured is the average tangential shear $\langle\gamma_{t}\rangle$ as a function of angular separation $\theta$.
@@
 (2023) for a detailed description of all the relevant systematic corrections that are required to yield a robust measurement of $\Delta\Sigma$.
+Amon et al.
+(2023) create the DES Y3+KiDS-1000 combined data vector used in this analysis by taking the inverse-variance weighted average of GGL measurements made using DES Y3 and KiDS-1000 separately.
+They show that these measurements can be combined because the KiDS-BOSS and DES-BOSS on-sky footprints have no overlap (see Figure 1) and they are statistically consistent (see Section 5.1.5 of Amon et al.
+(2023) for more detail).
+The error in the DES Y3 measurement is estimated using a Jackknife covariance.
+The lens galaxies and random points are split into 75 regions within the overlap area of 770 deg2 using the kmeans algorithm, resulting in each square region spanning 3 deg on a side or $\sim 65\;h^{-1}\rm Mpc$ (30 % larger than the largest angular scale measured).
+The KiDS-1000 error estimation is computed using the bootstrap method with square regions of 4 deg on each side equivalent to $\sim 80\;h^{-1}\;\rm Mpc$.
 2.2 kSZ Effect
@@
 The covariance matrix is constructed using the bootstrap method.
-[^2] Higher signal-to-noise kSZ measurements have recently been reported in (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a) using ACT DR6 data; our method could be extended in future to incorporate these data.
+[^2] Higher signal-to-noise kSZ measurements have recently been reported in (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b) using ACT DR6 data; our method could be extended in the future to incorporate these data.
 The coadded CMB temperature maps were created by combining the ACT DR5 day & night maps (Fowler et al., 2007; Swetz et al., 2011; Thornton et al., 2016; Henderson et al., 2016) with data from the Planck satellite (Planck Collaboration et al., 2020a).
@@
 (2021) perform a bootstrap resampling of the individual galaxies to estimate the covariance of the stacked kSZ profiles.
+Specifically, Schaan et al.
+(2021) repeatedly draw individual galaxies from the galaxy catalogue to generate a resampled galaxy catalogue of the same size.
+The kSZ profile is then measured on 10,000 realizations of the resampled galaxy catalogue which enable an estimation of the covariance matrix.
 The measurements at smaller aperture sizes have a dominant source of noise coming from the detector and foregrounds while the measurements at larger aperture sizes are dominated by large-scale CMB fluctuations.
@@
 The first term refers to the 1-halo density profile which corresponds to density from the effective central halo, and the second term refers to the 2-halo density profile which corresponds to contributions in the density coming from the surrounding neighbouring halos (Asgari et al., 2023).
-We model the transition between the 1-halo and 2-halo terms as a sum.
-This simplistic treatment of the transition region should suffice for the current level of precision in our analysis.
+We model the transition between the 1-halo and 2-halo terms as a sum rather than taking the maximum of the terms at any given scale.
+This simplistic treatment of the transition region should suffice for the current level of precision in our analysis, but future work that improves the precision of this model should explore the impact of this modelling choice.
 In the subsequent subsections below we will explain these 1-halo and 2-halo terms in more detail.
-We use a fixed Planck flat $\Lambda$CDM cosmology with the following values for parameters: $\Omega_{\rm m}=0.3111$, $\Omega_{\rm b}=0.0490$, $h=0.6766$, $n_{s}=0.9665$, $\sigma_{8}=0.8102$ (Planck Collaboration et al., 2020b).
-We fix the redshift of our model to be $\langle z\rangle=0.55$, the mean redshift of our lens sample (BOSS CMASS).
-This choice does not impact the results of our analysis when compared to a marginalization over the redshift distribution of the lens sample as shown in Amodeo et al.
-(2021).
-We further verified this to be true while including GGL.
+[…]
 3.1 The One Halo Term
@@
 In other words, we do not model the gas and dark matter around satellites and centrals with a Halo-Occupation Distribution (HOD) framework.
-This model ignores the potential systematic effect of satellites or ‘miscentering’, where satellite galaxies offset the stacking procedures used on the central galaxies in both the kSZ effect measurements and the GGL measurements.
+Additionally, this model ignores the potential systematic effect of satellites or ‘miscentering’, where satellite galaxies offset the stacking procedures used on the central galaxies in both the kSZ effect measurements and the GGL measurements.
+The impacts of ignoring satellites on the kSZ and GGL signals have been shown in (Hadzhiyska et al., 2023; McCarthy et al., 2024; Siegel et al., 2025; Bigwood et al., 2025).
+For CMASS galaxies Siegel et al.
+(2025) showed that ignoring satellites shifts halo mass estimates at the $\sim 10\%$ level, which is not significant enough to undermine the findings of this work.
+McCarthy et al.
+(2024) and Bigwood et al.
+(2025) showed with hydro simulations that for a CMASS-like sample in hydro simulations satellite fractions can be $\sim 10-20\%$ and have a scale dependent influence on the amplitude of both the kSZ and GGL signals at the $\sim 20\%$ level.
 We also ignore an additional uncertainty on the reconstruction bias factor $r_{v}$ which is about $\sim 10\%$ for the reconstruction used in this analysis (Schaan et al., 2021).
-We note that modelling of more precise measurements should include this effect (Ried Guachalla et al., 2024; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a).
+We note that modelling of more precise measurements should include this effect (Ried Guachalla et al., 2024; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b).
+The central goal of this work is to demonstrate the complementarity of measurements of the kSZ effect and GGL.
+This goal does not require a level of precision that warrants the accounting of these additional systematic effects.
+For this reason, it is suitable to make these modelling simplifications in this first iteration of the model framework.
+In Sections 4.2 and 4.3 we quantify how much these simplifications are expected to impact our parameter inference and by extension our analysis results.
+Future work designed for precise cosmological inference should account for these effects.
 3.1.1 Dark Matter
@@
 By using $\Delta=200$, our definition of halo mass $M_{\rm halo}$ is the same as $M_{200}$.
-This concentration relation was determined from N-body simulations for halo masses $10^{11}-10^{15}\;\rm M_{\odot}$ and redshifts $20<z<$ 2 which matches our selected sample.
+This concentration relation was determined from N-body simulations for halo masses $10^{11}-10^{15}\;\rm M_{\odot}$ and redshifts $0<z<2$ which matches our selected sample.
 We use a comoving $\bar{\rho}_{\rm m}$ because the radius $r$ is in comoving units.
@@
 This effect could bias our results at the few percent level (Yuan et al., 2021; Amon et al., 2023) and is a necessary improvement for future work.
-We also choose to ignore the back-reaction of baryonic feedback on the dark matter distribution: a second-order effect which is not necessary to model for our current level of precision (van Daalen et al., 2011b; Springel et al., 2018).
+We also choose to ignore the back-reaction of baryonic feedback on the dark matter distribution: a second-order effect which is not necessary to model for our current level of precision (van Daalen et al., 2011a; Springel et al., 2018).
 Our model of the 1-halo term closely follows an approach from Cromer et al.
@@
 $\log_{10}(M_{\rm halo})$ Halo Mass $\mathcal{U}(12.5,14.0)$ $13.42^{+0.44}_{-0.27}$ $13.41^{+0.08}_{-0.07}$ $13.44^{+0.06}_{-0.05}$
-$x_{c}$ Core scale fraction $\mathcal{U}(0.0,1.0)$ $>0.48$ $>0.46$ $>0.49$
+$x_{c}$ Core scale fraction $\mathcal{U}(0.0,1.0)$ >0.48>0.48 >0.46>0.46 >0.49>0.49
 $\alpha$ Intermediate slope
-$\beta$ Outer slope $\mathcal{U}(1,5)$ $2.47^{+0.34}_{-0.46}$ $<3.51$ $2.45^{+0.30}_{-0.23}$
+$\beta$ Outer slope $\mathcal{U}(1,5)$ $2.47^{+0.34}_{-0.46}$ <3.51<3.51 $2.45^{+0.30}_{-0.23}$
 $\gamma$ Inner slope
-$A_{\rm 2h}$ Two halo term amplitude $\mathcal{U}(0.0,1.5)$ $>0.631$ $0.66\pm 0.16$ $0.64\pm 0.16$
+$A_{\rm 2h}$ Two halo term amplitude $\mathcal{U}(0.0,1.5)$ >0.631>0.631 $0.66\pm 0.16$ $0.64\pm 0.16$
 Table 1: Summary of the halo model parameters, their priors, and the 68% confidence intervals for the fit to only kSZ measurements, only GGL measurements, and joint kSZ and GGL data.
@@
 We choose $\rho_{0}$ such that the mass ratio of baryons to total matter enclosed in a sphere of radius $r_{\rm b}$ is equal to the cosmic baryon fraction
-[…]
+$f_{\rm b}=\frac{\int_{0}^{r_{\rm b}}4\pi r^{2}dr\;\rho^{\rm 1h}_{\rm b}(r)}{\int_{0}^{r_{\rm b}}4\pi r^{2}dr\;\rho^{\rm 1h}_{\rm m}(r)}\;.$ (20)
 Using this condition, we solve for the baryon density amplitude $\rho_{0}$ to get
@@
 $\rho^{\rm 2h}_{\rm m}(r)=\bar{\rho}_{\rm m}\xi^{\rm 2h}_{\rm hm}(r)A_{\rm 2h}\;,$ (22)
-where $A_{\rm 2h}$ is an additional nuisance parameter to account for any potential systematic modelling errors in the 2-halo term, $\xi^{\rm 2h}_{\rm hm}(r)$ is the 2-halo term of the halo-matter correlation function, related to the matter correlation function by linear halo bias $\xi^{\rm 2h}_{\rm hm}=b(M)\xi^{\rm 2h}_{\rm mm}$, and $\bar{\rho}_{\rm m}$ is the comoving average cosmic matter density.
-We expect $A_{\rm 2h}$ to be less than unity to match the low amplitude of the 2-halo GGL signal Leauthaud et al.
-(2017); Amon et al.
-(2023) compared to predictions from clustering in Planck $\Lambda$CDM.
-As can be seen in Figure 2, the fit to only the kSZ effect measurements prefers much larger values of $A_{\rm 2h}$ than unity which is not physically justifiable.
-We choose our upper prior bound for $A_{\rm 2h}$ to minimize the prior volume effect (e.g.
-Ried Guachalla et al., 2025b) and reduce potential bias on the physical parameters of interest.
-We note that $M$ denotes the mass of each halo.
-The correlation function is related to the power spectrum by the inverse Fourier transform
+[…]
 $\xi_{\rm mm}^{\rm 2h}(r)=\int\frac{k^{2}dk}{2\pi^{2}}P^{\rm 2h}_{\rm mm}(k)j_{0}(kr)\;.$ (23)
 The 2-halo term of the matter power spectrum $P^{\rm 2h}_{\rm mm}(k)$ is given by
-[…]
+$P^{\rm 2h}_{\rm mm}(k)=P^{\rm lin}(k)\left[\int dM^{\prime}\frac{dn}{dM^{\prime}}b(M^{\prime})\hat{\rho}_{\rm NFW}(k)\right]^{2}\;,$ (24)
 where $P^{\rm lin}(k)$ is the linear matter power spectrum computed using the core cosmology library (pyccl) python package Chisari et al.
-(2019b), $dn/dM$ is the halo mass function Tinker et al.
+(2019a), $dn/dM$ is the halo mass function Tinker et al.
 (2008), $b(M)$ is the linear halo bias Tinker et al.
@@
 The excess surface mass density $\Delta\Sigma(R)$ is a function of the comoving surface mass density $\Sigma(R)$ which we define as
-[…]
+$\Sigma(R)=\int_{\rm LoS}\rho_{\rm m}\left(\sqrt{\chi^{2}+R^{2}}\right)d\chi\;,$ (34)
 where $\chi$ is the comoving line-of-sight distance to a redshift $z$ and $R$ refers to the projected comoving radius in the plane of the sky.
@@
 where the first term corresponds to the average comoving surface mass density enclosed within $R$ and is given by
-[…]
+$\overline{\Sigma}(<R)=\frac{2}{R^{2}}\int dR^{\prime}\;R^{\prime}\;\Sigma(R^{\prime})\;.$ (36)
 We compute the total matter excess surface mass density using the FFTLog algorithm implemented in pyccl Chisari et al.
-(2019b).
+(2019a).
 With this result we have a complete consistent framework for modelling the both the kSZ effect and GGL measurements.
@@
 The black points correspond to observations, the red corresponds to our fiducial model being fit to only the kSZ measurements, the blue corresponds to our fiducial model being fit to only the GGL measurements, and the purple corresponds to our joint fits to both the kSZ and GGL data.
-Dashed curves refer to the median (50th percentile) model of the posterior distribution and the bands denote the ±$\pm 1\sigma$ (32nd-68th percentiles) of the distribution of model curves.
+Dashed curves refer to the median (50th percentile) model of the posterior distribution and the bands denote the $\pm 1\sigma$ (32nd-68th percentiles) of the distribution of model curves.
 Lastly, solid curves denote the best-fit to the model.
@@
 (2013).
-To ensure convergence we ran the chains sufficiently long such that the Gelman-Rubin convergence parameter $\mathcal{R}$ was less than 1.01 for all model parameters Gelman & Rubin (1992).
+To ensure convergence we ran the chains sufficiently long such that the Gelman-Rubin convergence parameter $\mathcal{R}$ was less than 1.01 for all model parameters Gelman and Rubin (1992).
 4.1 The complementarity of kSZ and GGL
@@
 Figure 3 shows the posteriors for our model parameters obtained from fitting to the stacked kSZ profile data (red).
-The best-fit (minimum $\chi^{2}$) for this model is $\chi^{2}_{\rm min}/\mathrm{d.o.f}=21.8/14$ (PTE $=0.08$).
+The best-fit (minimum $\chi^{2}$) for this model is $\chi^{2}_{\rm min}/\mathrm{d.o.f}=21.8/16$ (PTE =0.15=0.15).
 As one would expect, the halo mass and amplitude of the 2-halo term are largely unconstrained when the model only has access to the kSZ profile.
@@
 The GGL measurements offer a lot of constraining power on $M_{\rm halo}$ and $A_{\rm 2h}$ but don’t have much constraining power for $\beta$ or $x_{\rm c}$.
-This fit to the data gave a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=2.5/7$ (PTE $=0.92$) implying that the GGL measurements were over-fit by our model.
+This fit to the data gave a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=2.5/9$ (PTE =0.98=0.98) implying that the GGL measurements were over-fit by our model.
 Furthermore, the GGL measurements are a probe of the total matter distribution which is comprised of mostly dark matter making it difficult to infer substantial information of the baryon density profile.
 The joint fit to the data is shown in purple.
-This fit yielded a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=24.8/25$ (PTE $=0.47$) implying a good fit to the data.
+This fit yielded a best-fit $\chi^{2}_{\rm min}/\mathrm{d.o.f}=24.8/26$ (PTE =0.53=0.53) implying a good fit to the data.
 In these purple contours we can see the overlapping regions of unconstrained red and blue contours leading to relatively tight constraints on all parameters except $x_{\rm c}$.
@@
 That is, a joint analysis of GGL and kSZ can account for the gas parameters, shrinking the width of the posterior back down, while the halo mass estimate undergoes a $\sim 20\%$ shift compared to the no-baryon model.
-We note that our joint kSZ+GGL constraints on $M_{\rm halo}$ are expected to be underestimated by $\lesssim 10\%$ compared to the mean halo mass of the sample $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$ likely due to the exclusion of miscentering in our model (Skibba et al., 2011; Hoshino et al., 2015).
+Our GGL+Baryons and kSZ+GGL constraints on $M_{\rm halo}$ are consistent with the mean halo mass $\langle M_{\rm halo}\rangle=3\times 10^{13}\;\rm M_{\odot}$ obtained with the SHMR Schaan et al.
+(2021) at the $1\sigma$ level but our GGL with no baryons constraint on halo mass shifts downward to agree at the $2\sigma$ level.
+We note that our mass estimate could be underestimated by $\lesssim 10\%$ due to the exclusion of miscentering in our model (Skibba et al., 2011; Hoshino et al., 2015).
+In future work we will improve our model by including satellites and accounting for miscentering which will yield an unbiased estimate of the halo mass.
 A crucial step in using kSZ profile measurements as a benchmark for feedback models and simulations is an accurate constraint on the halo mass.
@@
 (2024) had to use a large prior for $M_{\rm halo}:\mathcal{U}[5\times 10^{12},7\times 10^{13}]$.
-Having such a large uncertainty on halo masses make it difficult to interpret the strength of baryonic feedback from the stacked kSZ profiles measured by Schaan et al.
+Having such a large uncertainty on halo masses makes it difficult to interpret the strength of baryonic feedback from the stacked kSZ profiles measured by Schaan et al.
 (2021); Hadzhiyska et al.
@@
 (2024) are $\log_{10}(M_{\rm 500c}/\rm M_{\odot})=13.34\pm 0.04$.
-We find that our estimate of the halo mass is $\sim 30\%$ lower than what was found by McCarthy et al.
-(2024).
-Moreover, getting an unbiased estimate of the halo mass is crucial for modelling the tSZ effect (Moser et al., 2021, 2023) and the matter power spectrum suppression due to baryons (Debackere et al., 2020).
-Unlike the kSZ effect which scales linearly with halo mass $T_{\rm kSZ}(\theta)\propto M_{\rm halo}$, the tSZ effect has a much stronger scaling with halo mass $T_{\rm tSZ}(\theta)\propto M_{\rm halo}^{5/3}$.
-For this reason, an estimated halo mass that is biased high by $30\%$ will bias the model to over predict the amplitude of $T_{\rm tSZ}(\theta)$ by at least $50\%$.
-This potential bias arising from incorrect halo mass estimates could help alleviate discrepancies between the observed and predicted tSZ profiles in Moser et al.
-(2022) and tSZ clustering signals described in McCarthy et al.
-(2024).
-Similarly, a bias in halo mass estimates of $\sim 30\%$ could lead to an underestimation of the suppression of the matter power spectrum by $\sim 4\%$ at $k=1\;h/\mathrm{Mpc}$ (Debackere et al., 2020).
-By virtue of having an analytic model that can consistently model both the kSZ profile measurements and the GGL measurements we can completely side step any assumptions made to estimate the halo mass of the sample.
-In future work we will improve our model by including satellites and accounting for miscentering which will yield an entirely unbiased estimate of the halo mass.
+We find that our estimate of the halo mass is slightly lower than what was found by McCarthy et al.
+(2024) but still statistically consistent at the $2\sigma$ level.
+We note that there is additional uncertainty introduced by our method of converting between mass definitions due to the fact that the baryons are not described by an NFW profile, but the baryon fraction at $r_{\rm 500c}$ is lower than the cosmic baryon fraction and should only introduce uncertainty at the few percent level.
 4.3 Baryonic Feedback
 The kSZ effect probes the distribution of gas in the outskirts of halos.
-Only recently have direct measurements of the stacked kSZ effect profile had high enough SNR to be a powerful probe of the gas (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a).
+Only recently have direct measurements of the stacked kSZ effect profile had high enough SNR to be a powerful probe of the gas (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b).
 The ACT-BOSS kSZ measurements used in this analysis from Schaan et al.
@@
 The $\beta(M,z)$ relation was derived from gas profiles of resolved halos with masses of $M_{500}=10^{14}-10^{15}\;\rm M_{\odot}$, for this reason we acknowledge that our comparisons are to an extrapolation of the $\beta(M,z)$ relation of Battaglia (2016).
-We show our ±$\pm 1\sigma$ estimates on $\beta$ and $M_{\rm halo}$ compared to 2 models from Battaglia (2016).
+We show our $\pm 1\sigma$ estimates on $\beta$ and $M_{\rm halo}$ compared to 2 models from Battaglia (2016).
 The blue curve in Figure 6 is the Adiabatic model which corresponds to a simulation with only gravitational heating, the red curve is the AGN model which includes radiative cooling, star formation, supernova feedback, cosmic rays, and AGN feedback (more extreme feedback).
@@
 In Figure 4 we plot the average of the $11\leq\log_{10}(M_{*}/\rm M_{\odot})\leq 11.5$ bin and the $11.5\leq\log_{10}(M_{*}/\rm M_{\odot})\leq 12.0$ to most closely represent the expected suppression of our sample.
-[…]
+The difference in the suppression of $\Delta\Sigma$ between the older Illustris simulation and the TNG-300 simulation can be explained by the differences in subgrid model implementations for feedback.
+The Illustris model is known to have more extreme feedback than the TNG-300 model but is also known to have shortcomings in predicting observations such as gas fractions, galaxy color and morphologies, and several cluster properties (Nelson et al., 2019; Kauffmann et al., 2019; Rodriguez-Gomez et al., 2019).
+Our kSZ+GGL best-fit suggests a suppression of $\Delta\Sigma$ that gradually increases from 0% at 10 Mpc/$h$ to $15\%$ for $R\lesssim 1\;\mathrm{Mpc}/h$ (in comoving units) which appears to more closely reflect the Illustris feedback model than the IllustrisTNG-300 feedback model.
+Newer measurements of the kSZ effect profile in DESI-Y1 data also appear to have a shape that agrees with the older Illustris feedback model more than the IllustrisTNG-300 feedback model (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b).
 Our last point of comparison to simulations is directly to the baryon (gas) density profile shown in the leftmost panel of Figure 4.
@@
 We note that our model neglects satellites, for this reason we are neglecting the potential systematic of ’miscentering’.
-We expect that not accounting for this effect causes an underestimation of the kSZ amplitude in simulations by up to $\sim 20\%$ for CMASS-like galaxies ($12.111.5<\log M_{*}<$ 12.1) as shown in Figure 6 of Hadzhiyska et al.
+We expect that not accounting for this effect causes an underestimation of the kSZ amplitude in simulations by up to $\sim 20\%$ for CMASS-like galaxies ($11.5<\log M_{*}<12.1$) as shown in Figure 6 of Hadzhiyska et al.
 (2023), at the current level of statistical power, this effect would not significantly change the conclusions of this work.
-We note that this effect is dependent on the satellite fraction of the sample which means that miscentering is more significant for lower mass galaxies ($11.510.9<\log M_{*}<$ 11.5) and less significant for more massive galaxies ($12.712.1<\log M_{*}<$ 12.7).
+We note that this effect is dependent on the satellite fraction of the sample which means that miscentering is more significant for lower mass galaxies ($10.9<\log M_{*}<11.5$) and less significant for more massive galaxies ($12.1<\log M_{*}<12.7$).
 Considering Figure 2, we can see that ignoring miscentering from satellites in our model would most likely mean our values for $\beta$ are overestimated by $\sim 0.2-0.4$ which would not significantly impact the conclusions from this study.
@@
 In our model we use a fiducial value for the baryon radius $r_{\rm b}=10\times r_{200}$ (purple), but we include an additional data point with $r_{\rm b}=20\times r_{200}$ to show the range of values $\beta$ could be with this model depending on the choice of $r_{\rm b}$.
+All fits plotted from this work allow $x_{\rm c}$ to be a free parameter.
 5 Conclusions
@@
   • I.
-[…]
+    the kSZ effect serves as a complementary probe to GGL.
+By doing a joint fit to kSZ+GGL data, we can alleviate degeneracies in the model that exist when fitting to a single dataset (kSZ or GGL) as seen in Figures 2 and 3.
+When our fiducial model is fit to both kSZ and GGL measurements, the baryon density profile is tightly constrained at all relevant scales $r\approx 0.3-50$ Mpc/$h$ (in comoving units) as seen in the left panel of Figure 4.
+By having a tight constraint on the baryon density profile around the outskirts of the galactic halo, we determine empirically how baryons suppress $\Delta\Sigma$ at small scales as seen in the right panel of Figure 4.
+This work opens an avenue for using small scales in future cosmological weak lensing analyses.
+Additionally, not including baryons in a model of GGL can shift halo mass estimates by $\sim 20\%$ compared to a model that includes baryons fit to kSZ+GGL measurements as seen in Figure 5.
   • II.
@@
 Going forward, there are many avenues to build upon this analysis.
-First, new measurements of the stacked kSZ profile from ACT-DR6 and DESI-Y1 data using photometric (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2024; Hadzhiyska et al., 2024b) and spectroscopic (Ried Guachalla et al., 2025a) redshifts have significantly higher SNR than the measurements from Schaan et al.
+First, new measurements of the stacked kSZ profile from ACT-DR6 and DESI-Y1 data using photometric (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2024; Hadzhiyska et al., 2024b) and spectroscopic (Ried Guachalla et al., 2025b) redshifts have significantly higher SNR than the measurements from Schaan et al.
 (2021) (analysed in this work) with the ability to measure the signal in a more narrow range of redshift and mass.
@@
 [^7] https://github.com/aamon/Galaxy-galaxy-lensing-KiDS-DES-HSC Our modelling code glasz is built upon the pyccl Chisari et al.
-(2019b) and Mop-c-GT Amodeo et al.
+(2019a) and Mop-c-GT Amodeo et al.
 (2021) packages and is publicly available at the glasz homepage.
@@
 Our implementation of this model in glasz uses pyccl for all halo model computations rather than combining multiple cosmological codes.
+Appendix C Degeneracy in Baryon Profile Parameters
+The 2 baryon density profile parameters $x_{\rm c}$ and $\beta$ are highly degenerate with each other.
+While we have defined the core scale fraction $x_{\rm c}$ to span $[0,1]$ based off of the definition proposed previously in precursory work (Battaglia, 2016; Amodeo et al., 2021), this is simply a choice in the model.
+In Smith et al.
+(2018) the core scale fraction is fixed to $x_{\rm c}=2$ rather than 0.5.
+A higher value of $x_{\rm c}$ implies that the turnover in the density profile occurs further out into the halo which requires a higher (steeper) value $\beta$ for the baryon density to remain equivalent in the outskirts of the halo.
+To better understand this degeneracy between $x_{\rm c}$ and $\beta$, we fit the data with expanded priors on the $x_{\rm c}$ parameter.
+We find that only one of the two parameters can be constrained at a time.
+More specifically, when the MCMC chains reach a prior wall for one of the two parameters, the other parameter will be constrained.
+If the upper bound on the prior for $x_{\rm c}$ is 5, then $\beta$ will remain unconstrained.
+If the upper bound on the prior for $\beta$ is expanded to 10 in this case, then $\beta$ will be constrained while $x_{\rm c}$ remains unconstrained (the posteriors look similar to Figure 3 only shifted to higher values).
+Fixing $x_{\rm c}=2.0$ necessarily shifts the marginalized constraint on the outer power law slope to $\beta=3.60\pm 0.19$ while leaving $A_{\rm 2h}$ and $M_{\rm halo}$ unchanged.
+Ultimately, in order to make a fair comparison to the simulations of Battaglia (2016), we choose to keep the fiducial prior $\mathcal{U}(0.0,1.0)$ for $x_{\rm c}$ which is consistent with the model definition.